### v1

# Self-Supervised Skeleton Action Representation Learning: A Benchmark and Beyond

Abstract
Self-supervised learning (SSL), which aims to learn meaningful prior representations from unlabeled data, has been proven effective for label-efficient skeleton-based action understanding. Different from the image domain, skeleton data possesses sparser spatial structures and diverse representation forms, with the absence of background clues and the additional temporal dimension. This presents the new challenges for the pretext task design of spatial-temporal motion representation learning. Recently, many endeavors have been made for skeleton-based SSL and remarkable progress has been achieved. However, a systematic and thorough review is still lacking. In this paper, we conduct, for the first time, a comprehensive survey on self-supervised skeleton-based action representation learning, where various literature is organized according to their pre-training pretext task methodologies. Following the taxonomy of context-based, generative learning, and contrastive learning approaches, we make a thorough review and benchmark of existing works and shed light on the future possible directions. Our investigation demonstrates that most SSL works rely on the single paradigm, learning representations of a single level, and are evaluated on the action recognition task solely, which leaves the generalization power of skeleton SSL models under-explored. To this end, a novel and effective SSL method for skeleton is further proposed, which integrates multiple pretext tasks to jointly learn versatile representations of different granularity, substantially boosting the generalization capacity for different downstream tasks. Extensive experiments under three large-scale datasets demonstrate that the proposed method achieves the superior generalization performance on various downstream tasks, including recognition, retrieval, detection, and few-shot learning.
Index Terms:
Self-supervised learning, skeleton-based action understanding, contrastive learning, masked skeleton modeling
1 Introduction
Human activity understanding is an essential topic in the research of computer vision due to its wide applications in real life, such as human-robotics interaction [1], autonomous driving [2], and healthcare [3]. Among the different data modalities for actions, skeletons represent the human body by 3D coordinates of key body joints, which are lightweight, compact, and more robust to changes of view and background. Owing to these desirable advantages, skeleton has attracted much attention in human action analysis.
In the early works, many endeavors have been put into the supervised skeleton-based human activity understanding, e.g., recognition and detection [4, 5, 6]. However, these supervised methods heavily rely on huge amounts of labeled data, which requires time-consuming and expensive data annotation work, limiting the wide applications in the real world. As a remedy to this problem, self-supervised learning (SSL) attracts much attention and has been proven successful for representation learning. It exploits supervisory signals from unlabeled data, learning meaningful prior features and boosting generalization capacity of model for downstream tasks. Motivated by recent success in the image domain, great interest has arisen in adopting SSL for skeleton. However, it is not trivial to transfer these approaches into the skeleton data directly, which are with a more compact spatial structure, additional temporal dimension, and the absence of the background clues. To this end, researchers have made valuable exploration for skeleton-based SSL.
Generally, existing skeleton-based SSL works can be categorized into three types according to the pre-training pretext tasks, i.e., context-based, generative learning, and contrastive learning methods. The context-based methods construct the pseudo-label based on the intrinsic property of data, e.g., the joint angle prediction, to learn the spatial and temporal relations. Generative learning mainly focuses on reconstructing and predicting the skeleton data or the corresponding features. Contrastive learning methods model the high-level representations with an instance discrimination task. The various positive and negative skeleton views are generated by the well-designed spatial-temporal augmentations, boosting the consistency learning of the model.
Despite the huge progress recently, there is still a lack of a thorough literature review and analysis. Therefore, in this paper, we contribute a comprehensive survey of the self-supervised skeleton-based action representation learning. In contrast to other SSL surveys towards image, video, or text data, we focus on skeleton-based representation learning, which is the first literature to the best of our knowledge. This survey conducts a thorough review of mainstream SSL literature for skeleton and also involves the skeleton data collection, benchmark of performance, and the discussion of future possible directions. We believe our extensive work can benefit the research community and bring rich insights for future work.
Based on our literature review, it is noticed that most SSL methods for skeleton focus on the single paradigm, learning representations of single granularity, e.g., joint-level features (by masked skeleton modeling, MSM) [7, 8] or sequence-level features (by contrastive learning) [9, 10, 11, 12]. This limits the generalization capacity of the model to more downstream tasks. Although some works [13, 14, 15] make valuable efforts to combine different paradigms, they only achieve mediocre improvement due to the inherent gap of feature modeling mechanisms between the contrastive learning and masked modeling [16, 17]. To this end, a novel SSL approach for skeleton is proposed to fully boost the generalization capacity of SSL model, which integrates the contrastive learning and MSM to learn the joint, clip, and sequence level representations jointly. Specifically, we fully utilize the novel motion pattern exposed by the manual designed augmentations and model training for sequence-level contrastive representation learning, while adopting MSM for the joint-level feature modeling. Besides, we further propose a novel clip-level contrastive learning method, which significantly boosts the short-term model capacity, along with an effective post-distillation strategy to achieve a more compact representation space. Finally, extensive experiments under five downstream tasks, not limited to the single action recognition task used in most previous works, demonstrate the promising generalization capacity of the proposed method.
Our contributions can be summarized as follows:
  • To the best of our knowledge, we are the first to provide a thorough survey that comprehensively reviews the self-supervised skeleton action representation learning literature. Based on the taxonomy of context-based, generative learning, and contrastive learning, we give a detailed analysis of the pretext task design and highlight the special consideration for skeleton data along with the corresponding challenges.
  • We present a comprehensive benchmark of existing skeleton SSL works as well as a summary of the popular datasets and downstream tasks for SSL evaluation. Meanwhile, we provide insightful analysis from the perspective of model backbones and pre-training paradigms, and discuss the future possible directions.
  • Motivated by the limitations revealed by our survey, we explore skeleton-based versatile action representation learning to fully mine the generalization power of SSL models. An effective SSL schema is proposed, which integrates contrastive learning and MSM to jointly model the representations of different granularity, remarkably benefiting different downstream tasks.
  • We perform rigorous quantitative experiments to study the generalization efficacy of current skeleton-based SSL methods across five downstream tasks, including recognition, retrieval, detection, and few-shot learning, on both uncorrupted and corrupted skeletons. Under the proposed SSL method, the promising results are achieved for versatile action representation learning.
The remaining sections are organized as follows: We first present a thorough review in Sec. 2, for skeleton-based SSL representation learning. Subsequently, based on our investigation, we propose a new method exploring the combination of contrastive learning and masked modeling tasks in Sec. 3. Then, we comprehensively benchmark existing methods in Sec. 4, and verify the effectiveness of our proposed approach. Finally, we conclude and summarize with possible future directions in Sec. 5.
2 Review on Self-Supervised Skeleton-Based Action Representation Learning
Generally, a two-stage paradigm is utilized in skeleton-based SSL, i.e., pre-training on pretext tasks first and then fine-tuning on downstream tasks. In the pre-training stage, different pretext tasks are designed for deep neural networks, capturing training signals derived from the data itself, called the process of self-supervision [18]. After that, the learned knowledge as feature representations is transferred to downstream tasks as shown in Fig. 1 (c). In principle, this part is not only the goal of SSL, i.e., to improve downstream task performance with learned representations, but also the way to assess the quality of representation learning methods. Note that some downstream tasks, e.g., motion prediction and 2D-to-3D lifting, are not considered in this survey, because they essentially do not rely on the supervisory signal of human annotation and can serve as pre-training pretext tasks themselves, leading to possible unfair comparison. For reviews on these topics, we direct readers to [19, 20].
Next, we first introduce skeleton data as well as its collection in Sec. 2.1. Then, a review of the skeleton SSL methods is presented in Sec. 2.2 based on the taxonomy of pretext tasks methodologies as shown in Fig. 1 (b). A summary and discussion are finally provided in Sec. 2.3.
Figure 1: The taxonomy framework for self-supervised skeleton-based representation learning in our survey. The survey is structured around three dimensions: skeleton data collection, SSL pretext design, and SSL downstream task evaluation, providing a comprehensive review.
2.1 Skeleton Data as Human Representation
Skeleton represents a human motion as a sequence of the body keypoint coordinates over time. The collection methods of skeleton can be divided into two main categories, i.e., marker-based and markerless methods, shown in Fig. 1 (a).
Marker-based motion capture (Mocap) systems often rely on inertial measurement units (IMU) or reflective markers, placed on the body to track the movement of humans. It can provide reliable skeleton data with even sub-millimeter accuracy [21]. However, this method is too costly in many application scenarios and requires highly trained personnel to operate. Meanwhile, it suffers from great inconvenience, e.g., the time-consuming placement process and a requirement for a controlled environment. The datasets [22, 23] for generation tasks, e.g., pose estimation, are usually collected in this way to guarantee the accuracy.
Markerless Mocap often depends on deep learning algorithms for pose estimation from RGB and depth data. Video RGB data can be easily obtained from the Internet while we can also utilize the camera hardware, e.g., the depth camera including Microsoft Azure Kinect, and single or multiple RGB video cameras, to collect the RGB, infrared, and depth images in deployment. Then to get the 3D motion, computer vision algorithms, on multi-view geometry and pose estimation, i.e., OpenPose [24] are then employed to detect and extract joint center locations. The whole pipeline is presented in Fig. 1 (a). However, due to the limitations of hardware and estimation algorithms, such method can have large errors compared to marker-based methods [25]. Nevertheless, it is still chosen for most skeleton-based datasets [26, 27, 28] for its simpleness and convenience.
2.2 SSL Methods for Skeleton
With respect to the pretext tasks, most existing skeleton-based representation learning methods encompass three categories: (1) context-based, (2) generative learning, and (3) contrastive-learning methods. Based on the taxonomy in Fig. 2, we provide a comprehensive survey on the skeleton-based SSL works and highlight the special design for the skeleton as well as the corresponding challenges, to distinguish them from other data modalities.
2.2.1 Context-Based Methods
Context-based methods generate the supervisory training signals according to the inherent contextual information of provided data. The model is encouraged to learn the spatial-temporal relationships by training on the pre-defined task. Emerging from the image domain, the pretext tasks rely on the context understanding, e.g., rotation prediction [29]. In contrast, skeleton data introduces an additional temporal dimension, and possesses more compact spatial information, which presents a new challenge on how to mine the meaningful spatial-temporal context of skeleton by pretext tasks. Typically, there are three common types of context-based pretext tasks, view-invariance-based, temporal-order-based, and motion-prior-based method as shown in Fig. 3.
Figure 2: The taxonomy of the skeleton-based self-supervised learning methods in our review.
(a) View invariance learning.
(b) Temporal sorting.
(c) Motion prior knowledge learning.
Figure 3: Three types of context-based SSL methods for skeleton.
1) View Invariance. Due to the variance of the observation viewpoint, the skeleton estimation can suffer from occlusion and noise. Therefore, learning view-invariant representations should be beneficial for action recognition task, and has been widely studied in the supervised skeleton-based action recognition task [30]. Li et al[31] proposed a view classification pretext task for unsupervised action representation learning. Specifically, as shown in Fig. 3 (a), skeleton sequences of different views are fed into the encoder, subsequent to which a view classifier predicts the view labels. To learn the view-invariant features, a Gradient Reversal Layer [32] is added to reverse the optimization direction of the encoder. Resort to this, the encoder can learn the features insensitive to view in an adversarial manner. Likewise, Paoletti et al[33] adopted the rotation prediction, i.e., pitch angle, yaw angle, and roll angle, as well as a gradient reverse operation to achieve the viewpoint-invariance learning.
There are also some subsequent SSL works [34, 35, 36] focusing on learning view-invariant representations. However, they adopted a contrastive learning paradigm to explicitly learn the alignment between the two views and showed a significant performance improvement.
2) Temporal Sorting. In this pretext task, skeleton data is treated as temporal sequences, and shuffled randomly. The model takes the shuffled skeleton as input and outputs the corresponding shuffling pseudo label to restore the temporal order. Specifically, the skeleton sequence is usually divided into multiple clips first, each of which contains several consecutive frames. Then the shuffling operation takes place at the clip level rather than the frame level because it is difficult to capture motion patterns by the difference between two adjacent frames. The model is trained to predict the shuffling label as shown in Fig. 3 (b), e.g., with a Cross-Entropy loss.
Some skeleton-based SSL works [13, 37] are equipped with temporal sorting to model the temporal dependencies, similar to SSL works in video [38, 39]. However, such pretext task is often integrated with other pretext tasks, because it only models temporal features explicitly, leaving the crucial spatial representations of skeleton under-explored. Therefore, new pretext tasks are studied to jointly learn the spatial-temporal relationships of skeleton.
3) Motion Prior. The motion dynamics of skeleton joints contain rich semantic information beneficial for action understanding. Therefore, researchers propose to leverage the inherent motion prior knowledge to generate pseudo labels for pre-training as shown in Fig. 3 (c). Cheng et al[40] designed a movement direction prediction task. The model is constrained to estimate the direction of the instantaneous joint velocity, i.e., whether the targeted joints are moving in a positive direction. Based on this, Kim et al[41] proposed a multi-interval pose displacement strategy to encourage both local and global attention learning. The key idea is to predict the motion direction and magnitude of the central joint and other joints during a random interval. On the other hand, Yang et al[42] introduced more motion prior information as the pseudo label, i.e., intra-joint motion variance, inter-joint motion covariance, intra-frame joint angle, and inter-frame motion deviation. The model is pre-trained by a regression task to predict this prior knowledge.
In summary, compared with image and video modalities, the context-based pretext tasks for skeleton usually involve more spatial-temporal design and motion dynamics modeling. However, these tasks are often designed manually and hard to capture the high-level underlying semantic distribution. Therefore, more efforts have been paid to the generative and contrastive learning SSL methods, which will be discussed in the following.
2.2.2 Generative Learning Methods
Generative methods utilize the generative power of the neural network to capture the spatial-temporal co-occurrence relationships among skeleton joints, modeling the underlying data distributions. In this case, the meaningful representations are learned by reconstructing or predicting the related signals of input skeletons. Here we do not strictly distinguish between “reconstruction” and “prediction” for generation, but use “reconstruct” to refer uniformly. Generally, the goal of the generative learning tasks can be formulated as:
$\centering argmin_{\theta}\ \mathcal{L}(\mathcal{D}(\mathcal{E}(\mathcal{T}_{in}(x))),\mathcal{T}_{tar}(x);\theta),\@add@centering$ (1)
where $x$ is a skeleton sequence, $\mathcal{D}$ and $\mathcal{E}$ are the decoder and encoder, respectively. $\mathcal{T}_{in}$ is the transformation applied to the original data, e.g., random masking, while $\mathcal{T}_{tar}$ maps the input into the target space where the loss objective $\mathcal{L}$ is applied to optimize the model parameters $\theta$.
When $\mathcal{T}_{in}$ and $\mathcal{T}_{tar}$ are both the identity function, the model is constrained to reconstruct the input data simply, known as AutoEncoder (AE). It naturally creates an information bottleneck to achieve dimension reduction, mapping data from input space onto low-dimensional feature space. Due to its simpleness, AE is used to learn prior representations in earlier skeleton-based works [14, 43]. However, since the skeleton data is still redundant in spatial and temporal dimensions, the encoder tends to memorize the input data at a low level instead of modeling the high-level semantic knowledge. Then, denoising AE (DAE) is studied for alleviating this shortcut [44, 45]. For example, Nie et al[45] applied a series of view corruptions and constrained the model to reconstruct the clean data to disentangle the view and pose features. On the other hand, inspired by the remarkable success of masked language/image modeling (MLM, MIM), which can be regarded as a variant of DAE, researchers have aroused a surge of interest in exploring masked skeleton modeling (MSM) for human action representation learning. The earlier works are mainly based on recurrent neural network (RNN) or graph convolutional network (GCN) to predict the masked skeletons. However, lacking good scalability, these models only show mediocre performance improvement. Recently, inspired by the success of Vision Transformer (ViT) as masked autoencoders (MAE) [46], Transformer-based models have been explored for the masked skeleton modeling [47, 48]. However, due to the relative redundancy in skeleton and the lack of a large-scale dataset as ImageNet, directly applying Transformer for masked skeleton reconstruction can suffer from the over-fitting problem. To this end, researchers have made endeavors on two crucial designs, masking strategy and reconstruction target, which are discussed as follows.
1) Masking Strategy has been proved crucial as the design of $\mathcal{T}_{in}$ in Eq. (1) for MIM. In the field of RGB images, MAE [46] adopts a patch-based masking strategy with a large masking ratio of 75%. For the skeleton data, researchers have fully explored the masking strategy in the spatial-temporal dimension. Although the optimal masking strategy and ratio can differ with the encoder backbone and MSM task setting, most works have found that spatially body-part-based and temporally segment-based strategy with large masking ratio, e.g., 90% in [10], can produce decent results. Concretely, body-part level masking refers to performing masking regarding the different joints in a body part as a whole, e.g., torso and left leg, while the temporal segment-based strategy masks the same joints across consecutive frames. These designs aim to reduce the shortcuts caused by spatial-temporal redundancy in skeleton, which is in line with the observation in images [49] and videos [50].
Recently, more elaborate masking strategies have been studied to achieve more effective exploitation of valuable semantic information. The work [51] found that masking the limbs is always better than masking the torso and head, especially the right hand and leg. MAMP [48] proposes a motion-aware masking strategy. The moving parts are located and masked by calculating the motion displacement between adjacent frames, which achieves better results than the random masking strategy. These results demonstrate that masking motion regions, which are often semantic-rich, promotes the model to learn more meaningful features in masked motion modeling. It can be also explained by the theory in [52], i.e., because samples from the same action often contain similar motion patterns, this masking strategy on the motion regions implicitly leads to better alignment of mask-induced positive pairs, achieving more discriminative feature spaces.
Meanwhile, some literature adopts special mask designs, yielding new MSM task forms. MS2[13] directly utilizes temporal frame-level masks to train the model on the motion prediction task. MotionBERT [53] adopts a 2D-to-3D lifting pretext task, in which it masks all the values in depth $z$ channel and partially in $x$ and $y$ channels, encouraging the model to predict the original 3D skeletons. These works can be divided into masked skeleton modeling in a broad sense.
2) Reconstruction Target, known as the implementation of $\mathcal{T}_{tar}$, also varies from the input space to the feature space. Most existing works reconstruct the skeleton data in the input coordinate space using MSE loss. Some works [15] propose to perform temporally reverse reconstruction to learn the skeleton dynamics instead of trivial representations that just remember the input.
Instead of directly reconstructing the input, Mao et al[48] proposed to reconstruct the motion in MSM, i.e., the difference of the corresponding joints between adjacent frames. The work [54] formulates the skeleton as an unordered 3D point cloud and maps the 3D data onto color space. This mapping function is artificially defined according to the spatial-temporal relationship of skeleton joints. Therefore, the model can learn the spatial relation and temporal dependency by reconstructing the color of the skeleton cloud.
2.2.3 Contrastive-Learning-Based Methods
Contrastive learning has been proven effective for different data modalities, e.g., images, point cloud, as well as skeleton data. Generally, contrastive learning pursues the consistency of the positive samples, which are usually the augmented counterparts of the original data. MoCo [49] utilizes the negative samples to establish an instance discrimination pretext task. Meanwhile, self-distillation [55] and feature decoupling [56, 57] concepts have also been explored in contrastive learning. These early pioneering works [49, 58, 55, 56], have made a huge impact and encouraged unique designs for skeleton contrastive learning.
For skeleton contrastive representation learning, most existing methods are based on MoCo v2 [49]. It adopts an asynchronously momentum-updated key encoder and an online query encoder, along with a memory queue to store a large number of consistent negative samples. Specifically, different augmentations are applied to the skeleton $x$ to generate the positive pair $(x_{q},x_{k})$, while the negative sample features $m_{i}$ are stored in a memory queue. The model is constrained to retrieve the positives among the negative samples, optimizing the following InfoNCE objective:
$\mathcal{L}_{Info}(z_{q},z_{k})=-\log\frac{\exp(z_{q}\cdot z_{k}/\tau)}{\exp(z_{q}\cdot z_{k}/\tau)+\sum_{i=1}{\exp(z_{q}\cdot m_{i}/\tau)}},$ (2)
where the $z_{q}$/$z_{k}$ is the query/key embedding encoded by the query/key model and $\tau$ is the temperature hyper-parameter. $m_{i}$ is the $i$-th feature anchor as the negative sample. Based on this, we review current skeleton contrastive learning methods from three aspects of design, i.e., augmentation design and learning strategy, objective engineering, and cross-modal learning.
1) Augmentation Design and Learning Strategy. Data augmentation exposes novel motion patterns and generates diverse positive views, which have been found crucial to the success of contrastive learning [59, 60]. Different from images, data augmentation for skeleton is relatively less developed. Therefore, the earlier works mainly focus on exploring practical spatial-temporal augmentations for skeleton contrastive learning [61, 62, 63]. For example, Rao et al[61] proposed a series of augmentations including Rotation, Shear, Reverse, Gaussian Noise, Gaussian Blur, Joint Mask and Channel Mask, some of which are used as the default basic augmentations in the future research. Notably, the optimal augmentations for different backbones are usually different. For example, the Joint Mask is found detrimental for GCNs [12], while beneficial for GRU model [62], which implies that GCNs are more sensitive to the spatial corruption. Generally, this difference arises on account of different modeling mechanisms and model capacities, which further increases the difficulty of the augmentation design and selection.
Inspired by the success of mixing-based augmentations for images, Chen et al[64] proposed SkeleMixCLR equipped with SkeleMix augmentation which combines the topological information of different skeleton sequences. To learn from this mixed skeleton, SkeleMixCLR obtains the corresponding part level and the whole-body level features and pursues the local-global invariance. Based on this, SkeAttnCLR [65] further designs an attention mechanism to perform local contrastive learning on salient and non-salient features. Due to the generated novel input views and regularization effect on the feature space, mixing augmentation often leads to consistent improvement for representations.
To further improve consistency learning, researchers have made endeavors to introduce more and stronger data augmentations. However, the ensuing problem is the over-distortion [66] of the augmented data, leading to the model performance degradation. In other words, some strong augmentations would seriously corrupt the semantic information, change the data distribution, and result in the difficulty of model consistency learning. To address this, AimCLR [60] utilizes two branches to encode the weakly and strongly augmented views, respectively. Then the model optimizes the similarity of the distributions output by two branches as a soft consistency learning target:
$\mathcal{L}_{Soft}$ $=KL(p(z|z_{weak})\|p(z|z_{strong})),$ (3)
${p}\left(z|z_{*}\right)$ $=\frac{\exp(z\cdot z_{*}/\tau)}{\exp(z_{k}\cdot z_{*}/\tau)+\sum_{i=1}^{M}{\exp(m_{i}\cdot z_{*}/\tau)}},$
where $z_{weak}$ and $z_{strong}$ are the corresponding embeddings of weakly and strongly augmented views. Instead of the one-hot target, this objective utilizes the similarity distribution as the soft target to guide consistency learning, which can be viewed as a self-distillation process. Further, Zhang et al[12] introduced more strong augmentations, e.g., randomly dropping the skeleton edges and joints, and proposed a hierarchical contrastive learning framework. It performs a decoupled progressive augmentation invariance learning by optimizing the consistency only between the augmented samples with adjacent strength, where the weakly augmented branch serves as the mimic target of the strongly augmented branch. These works show that the weakly augmented views can effectively guide the learning of the corresponding strongly augmented samples, leading to more stable representation learning and improvement.
In addition, Lin et al[67] explicitly distinguished the static regions and motion regions, namely, actionlet, in human skeleton, and introduced a motion-aware augmentation strategy. By mining the actionlet in an unsupervised manner, the semantic-reserving augmentations are employed for actionlet regions, while the noise perturbations for non-actionlet regions, avoiding the over-distortion problem. On the other hand, instead of augmenting at the input level, HaLP [68] proposes a latent positive hallucinating method by exploring the latent space around the corresponding prototype.
In summary, the development of data augmentations along with the corresponding learning strategy significantly boosts the performance of skeleton contrastive learning, which has always been an important and popular topic for 3D skeleton contrastive learning.
2) Objective Engineering. In addition to the widely used InfoNCE loss, some SSL works have also made efforts to explore new loss functions for extra regularization or new objectives. Here we introduce them from the perspective of False Negative (FN) problem, which widely exists in skeleton contrastive learning based on negative examples, e.g., MoCo v2. Concretely, false negatives refer to the negative samples but from the same semantic category. Traditional contrastive learning relies on the one-hot label and directly pushes them away in the feature space, which forces the model to discard the shared semantic information and leads to slow convergence [69]. We point out that, due to the lack of description of objects and backgrounds, there are fewer action categories represented by skeleton data, which leads to a more serious FN problem. To this end, the following aspects are considered to tackle this issue.
The first straightforward solution is based on the false negative detection and cancellation [9, 60, 70]. These methods first calculate the similarity between the sample and negatives, selecting the top $k$ negatives that are most similar as the potential false negatives, which are then involved as extended positives in contrastive learning.
Some other methods [10, 71] suggest employing the distillation objective as an adaptive re-weighting regularization for the one-hot instance discrimination pretext task, which can be formulated as follows:
$\mathcal{L}_{KD}$ $=-{p}\left({z}_{k},\tau_{k}\right)\log{p}\left(z_{q},\tau_{q}\right),$ (4)
${p_{j}}\left({z},\tau\right)$ $=\frac{\exp({z}\cdot{m}_{j}/\tau)}{\sum_{i=1}{\exp({z}\cdot{m}_{i}/\tau)}}.$
They assign the attraction weights to negatives based on the calculated similarity. Specifically, if a sample possesses high similarity with the positive, a larger attraction weight would be assigned to involve it in similarity optimization.
Besides, prototype-based contrastive learning is also studied for high-level semantic consistency learning. It usually performs clustering in feature space to assign instances to different cluster prototypes as a pseudo-semantic label. Then the model learns more high-level semantics by contrasting different prototypes. By virtue of such a way, the model pays more attention to the cluster-level discrimination task rather than the instance-level, which alleviates the false negative problem.
In addition to the above designs, as introduced above, some negative-sample-free contrastive frameworks have been developed recently, which avoid the difficulty of explicitly specifying negative examples. Based on BYOL [55], the works [72, 73] performs the positive-only consistency learning in the hyperbolic space. Some other works [74, 75], inspired by the feature decorrelation concept, propose to learn the decorrelated representations based on Barlow twins [56] and Variance-Invariance-Covariance Regularization (VICReg) [57] as the framework.
Figure 4: Different representations of skeleton data. From left to right are time series, 2D pseudo-image, spatial-temporal graph.
3) Cross-Modal Learning. Skeleton naturally provides different data modalities to represent human movement, e.g., motion and bone modality defined in the previous works [5, 76]. Meanwhile, many works [5, 76] have proved that the fusion of different modal knowledge can be beneficial to achieve richer and more representative action depictions. Therefore, cross-modal learning has attracted the attention of researchers for skeleton representation learning.
CRRL [15] performs contrastive learning between the joint and motion modalities, where the encoders for motion and joint data are homologous, i.e., obtained by momentum-updated strategy. However, this design can be unreasonable because the model can have difficulty dealing with two modalities simultaneously. In contrast, CrosSCLR [9] adopts three separate encoders for joint, bone, and motion data, and aims to align the distributions among the neighborhoods in latent space. Intuitively, it encourages the sample embedding together with its neighbors of modality $v$ should also be close in the latent space of modality $u$. However, this leads to a two-stage training process, as the model needs to be pre-trained on a single view first to obtain reliable unique modal knowledge. To this end, CMD [10] proposes a general bidirectional distillation objective, where the modal knowledge is directly modeled as the whole similarity distribution in the customized latent space. The proposed cross-modal distillation objective and the instance discrimination task jointly optimize the model, yielding a concise single training phase. On the other hand, to get rid of the limitation of using separate encoders for different modalities, UmURL [75] develops a unified modality-agnostic encoder, which can handle different modal inputs relying on the modality-specific input embedding layer and feature projection layer.
Meanwhile, we can also organize skeleton data into different representations as shown in Fig. 4, e.g., the temporal series, 2D pseudo-images with frame and joint dimensions, and spatial-temporal graphs with nodes and edges. Following this concept, ISC [62] encodes these different skeleton representations with different model backbones, which are then projected into a shared latent space. A cross-modal contrastive learning loss is applied for pre-training.
4) Others. In addition to the aforementioned literature, some works also employ contrastive learning. Considering the hierarchical structure of skeleton data, HiCo [11] performs contrastive learning in a hierarchical manner to model features of different levels. Some works [13, 37, 77, 15] combine the contrastive learning paradigm with others in a multi-tasking manner. Notably, in addition to simply combining reconstruction and contrastive learning pretext tasks, PCM3 [71] proposes a collaborative design to further improve the representation learning. Eq-Contrast [78] formulates the temporal sorting task into an equivariant contrastive learning objective for multi-task pre-training. Meanwhile, not limited to SSL, contrastive learning is also studied in supervised [79] and semi-supervised [80, 81] learning for skeleton.
2.3 Summary and Discussion
The skeleton-based SSL literature is categorized into context-based, generative learning, and contrastive learning types. All these methods aims to capture joint features as well as their relationships from unlabeled skeleton data, to obtain a meaningful representation of the motion. Due to the lack of ground-truth labels, these methodologies introduce prior knowledge in different means to enable representation learning.
Specifically, context-based methods introduce the prior knowledge by the predefined pretext-task designs, e.g., predicting the equivariant properties of some specific transformation, learning the spatial concurrence or temporal dependencies of skeleton joints. Generative approaches generally reconstruct or predict the skeleton target without much prior information. However, some recent works have shown that some special designs, e.g., masking the motion regions, can further boost the performance, where the prior knowledge is also utilized that the moving parts can better represent the action. In the context of contrastive learning, the prior information is mostly reflected through the data augmentations for the construction of positive pairs, which affect the invariance learning of transformation-related features. Overall, although downstream tasks are often unknowable in pre-training, these methods can obtain a potentially beneficial representation space by introducing different prior knowledge.
3 The Proposed Method: Prompted Contrast with Masked Motion Modeling
3.1 Motivation
By reviewing the previous works, we find that most existing skeleton-based SSL methods employ a single paradigm or simply combine different methods, evaluated on the action recognition task solely. We argue this ignores the hierarchical design for representation learning of different granularity and leaves the generalization capacity of model leashed, as we will verify later in Sec. 4.7. To construct a strong baseline model for skeleton-based SSL, we propose prompted contrast with masked motion modeling (PCM3++), which combines contrastive learning and MSM by exploring the potential collaboration. It enables versatile representation learning by joint, clip, and sequence level feature modeling, and significantly improves the performance of different downstream tasks.
Specifically, our method is based on our previous work [71] and integrates valuable improvement for better versatile representation learning. We first briefly introduce the joint-level and sequence-level feature modeling of the baseline in Sec. 3.2 - 3.4. Based on this, we introduce a novel asymmetric clip-level contrastive learning method in Sec. 3.2. The training strategy is presented in Sec. 3.5, including the newly proposed post-distillation policy for further representation refinement. The readers can refer to [71] for more details.
3.2 Skeleton Contrastive Learning
Our pipeline for contrastive learning follows MoCo V2 [49]. For a positive pair $(z,z^{\prime})$, the model optimizes the InfoNCE objective defined in Eq. (2) along with a distillation objective in Eq. (4) as regularization:
$\centering\mathcal{L}_{CL}(z,z^{\prime})=\mathcal{L}_{Info}(z,z^{\prime})+\mathcal{L}_{KD}(z,z^{\prime}).\@add@centering$ (5)
To boost the consistency learning, we propose a series of effective spatial-temporal augmentations to construct diverse positive pairs, which are introduced as follows.
1) Spatial Intra- and Inter- Skeleton Contrastive Learning. For intra-skeleton augmentations, we utilize Temporal Crop-Resize, Shear, and Joint Jittering to construct the positive pair $(s_{intra},s^{\prime})$. Then, we obtain the corresponding representations $(z_{intra},z^{\prime})$ via the query/key encoder $f_{q}(\cdot)/f_{k}(\cdot)$ and embedding projector $h_{q}(\cdot)/h_{k}(\cdot)$, respectively.
Meanwhile, Mixing augmentations are adopted to construct inter-transformed views. Specifically, given two skeleton sequences $s_{1},s_{2}$, we obtain the mixed skeleton data $s_{inter}$, e.g., by Mixup. Subsequently, we can obtain the embeddings corresponding to the mixed data by $z_{inter}=h_{q}\circ f_{q}(s_{inter})$. The optimized objective is for $(z_{inter},z_{inter}^{\prime})$, where $z_{inter}^{\prime}=(1-\lambda)(h_{k}\circ f_{k}(s_{1}))+\lambda(h_{k}\circ f_{k}(s_{2}))$.
2) Temporal Asymmetric Clip Contrastive Learning. The above designs focus on the sequence-level consistency modeling, i.e., the augmented data often contains sufficient temporal motion information. However, it ignores the short-term motion representation learning of model, which is necessitated for downstream tasks with dense prediction such as action detection. Meanwhile, the consistency learning for short clips, which can be viewed as a challenging temporal augmentation, can also benefit the whole representation quality learned by model. To this end, we present the clip-level skeleton contrastive learning design in this part.
To sample a motion sub-sequence (clip), we can apply the Temporal Crop transformation, and the obtained clip is much shorter ($10\%\sim 40\%$) than the original sequence. However, the clip only contains partial motion information and directly aligning the semantic embeddings of two clips is difficult for model, leading to the unstable training. Therefore, we introduce an asymmetric design, which only feeds the short clip into query branch, while the key branch takes the normal original sequence as an anchor. Meanwhile, we keep the two augmented data possessing the same spatial views to decouple the temporal variation. Specifically, we first apply the intra-augmentations to a sample $s$ to generate the $s_{ach}$. Then we further utilize the Temporal Crop to sample a continuous clip $s_{clip}$ from $s_{ach}$. It means that, $s_{ach}$ is subsuming $s_{clip}$ in temporal dimension while sharing exactly the same spatial transformations, to provide a precise and feasible target for clip-level consistency learning. Similarly, the model takes the positive pair $(z_{clip},z_{ach})$ for optimization, where $z_{clip}=h_{q}\circ f_{q}(s_{clip})$. In implementation, we just mask $z_{ach}=z^{\prime}$ for efficiency.
3.3 Masked Skeleton Prediction
To model the joint-level representations, we utilize the masked skeleton modeling task with a segment-wise masking strategy at body part level. To predict the masked regions from masked skeleton $s_{mask}$, we employ a decoder $dec(\cdot)$ taking the representations from the encoder as input. The MSE loss between original data $s$ and predicted data $s_{predict}$ by decoder is optimized in the masked regions:
$\mathcal{L}_{Mask}=\mathbb{E}\left(||(s-dec\circ f_{q}(s_{mask}))\odot(\mathbf{1}-M)||_{2}\right),$ (6)
where $M$ is the binary mask and $\mathbf{1}$ is an all-one matrix with the same shape as $M$.
3.4 On the Connection of Contrastive Learning and Masked Prediction
As discussed in the work [71], simply combining these two tasks can be sub-optimal due to the inherent gap of their feature modeling mechanisms [16, 17]. To this end, we explore the potential synergy for exploitation between them.
1) Novel Positive Pairs as Connection. We utilize special data views during the masked prediction training to provide more diverse positive samples for contrastive learning. First, the masked skeleton view $s_{mask}$ naturally simulates the occlusion for skeletons, serving as challenging positives. Meanwhile, we also take the predicted skeleton view $s_{predict}$ output by decoder $dec(\cdot)$ as positive samples. It contains the inherent uncertainty and diversity brought by continuous training of the model, which contributes to encoding more diverse movement patterns. On the other hand, the semantic consistency of the output skeleton with respect to the model itself is encouraged, i.e., the predicted view can also be perceived well by the encoder, connecting the low-level reconstruction with the high-level semantic modeling.
In a nutshell, we utilize the masked view $s_{mask}$ and the predicted view $s_{predict}$ as positives. Together with the manually constructed positive pairs in Sec. 3.2, we present all positive (embedding) pairs $\{(z_{q},z_{k})\}$ as follows:
$\{(z_{q},z_{k})\}=\{($ $z_{intra},z^{\prime}),(z_{inter},z^{\prime}_{inter}),(z_{clip},z^{\prime}),$ (7)
(((( $z_{mask},z^{\prime}),(z_{predict},z^{\prime})\}.$
2) High-Level Semantic Guidance. Note that we feed the predicted view into the contrastive learning pipeline. The gradients of $s_{predict}$ from the contrastive learning branch are propagated to update the reconstructed decoder $dec(\cdot)$. It provides the high-level semantic guidance for the skeleton prediction together with the joint-level MSE loss in Eq. (6), leading to better masked prediction learning and higher quality of $s_{predict}$ as positive samples.
3.5 The Whole Training Strategy
Overall, the model jointly optimizes the contrastive learning and masked skeleton modeling. Here we present the overall objective and the training strategy. For clarity, we utilize $\mathcal{L}_{Info}^{all}$ and $\mathcal{L}_{KD}^{all}$ to represent the sum of the losses for each positive pair defined in Eq. (7). It can be formulated as:
$\mathcal{L}_{Info}^{all}=\sum_{(z_{q},z_{k})}\mathcal{L}_{Info}^{query},\ \ \mathcal{L}_{KD}^{all}=\sum_{(z_{q},z_{k})}\mathcal{L}_{KD}^{query}.$
$\mathcal{L}_{Info}^{query}$ and $\mathcal{L}_{KD}^{query}$ are the respective terms for the specific $query$ view summarized in Eq. (7). During training, the following objective is applied to the whole model:
$\mathcal{L}=\mathcal{L}_{Info}^{all}+\lambda_{m}\mathcal{L}_{Mask}+\lambda_{kd}\mathcal{L}_{KD}^{all},$ (8)
where $\lambda_{m}$ and $\lambda_{kd}$ are the loss weight. Based on this, we further propose two advanced training strategies to boost the representation learning.
1) Prompted Multi-Task Pre-Training. Considering the difficulty of encoding different data views simultaneously, we employ prompt-based guidance to assist the model to learn from different data views explicitly. Specifically, we attend to a series of domain-specific prompts for different augmented views, i.e., $p_{intra},p_{inter},p_{clip},p_{mask}$, and $p_{predict}$. Their dimension equals to the spatial size of the skeleton data. Then, these domain-specific prompts are added to the corresponding skeleton ($s_{*}$ means * view of $s$):
$s_{*}=s_{*}+p_{*}.$ (9)
These decorated skeletons are fed into the query/key encoder for self-supervised pre-training, providing the training guidance and achieving better representations.
2) Post-Distillation Refinement. To further improve the representation quality, we introduce a post-distillation strategy as an optional refinement process. After obtaining a good prior feature space by Eq. (8), we directly remove the one-hot label constraint in InfoNCE objective $\mathcal{L}_{Info}^{all}$, and only apply the soft label, i.e., the distillation loss $\mathcal{L}_{KD}^{all}$, which assigns the attraction weights adaptively according to the similarity. This can be seen as a more explicit feature clustering process to obtain a more compact representation space by alleviating the false negative problem.
4 Benchmark Skeleton SSL Methods
We provide a comprehensive benchmark of existing methods, along with the skeleton datasets, backbone architectures, and downstream tasks in this section. Meanwhile, the evaluation of proposed PCM3++ is reported.
TABLE I: Summary of popular skeleton datasets. Marker denotes the dataset is collected by Marker-MoCap methods.
Dataset’Year Instance Class. Sub. View Joint Marker
CMU [82]’03 2235 45 144 - 31
SBU [83]’12 300 8 7 - 15
NW-UCLA [84]’14 1,494 10 10 3 20
UWA3D [85]’15 1,075 30 10 5 15
SYSU [86]’15 480 12 40 1 20
NTU 60 [26]’16 56,880 60 40 80 25
PKUMMD [28]’17 21,545 51 66 3 25
Kinetics [87]’17 $\sim$260k 400 - - 18
TSU [88]’19 16,115 31 18 - 17
NTU 120 [27]’19 114,480 120 106 155 25
UAV [89]’21 22,476 155 119 17
BABEL-60 [90]’21 45473 60 - - 25
BABEL-120 [90]’21 48,978 120 - - 25
4.1 Datasets
We first summarize the popular datasets for skeleton action understanding in Table I. Among them, the following datasets are widely used in skeleton-based SSL evaluation.
1) NTU RGB+D 60 Dataset (NTU 60) [26] is the most popular skeleton dataset. There are 56,578 videos with 25 joints for a human, captured by three Microsoft Kinect v2 cameras. The skeleton sequences are divided into 60 action categories, performed by 40 volunteers. Two evaluation protocols are recommended: a) Cross-Subject (xsub): the data for training are collected from 20 subjects, while the other 20 subjects are for testing. b) Cross-View (xview): the training set consists of front and two side views of the action performers, while testing set includes the left and right 45 degree views.
2) NTU RGB+D 120 Dataset (NTU 120) [27] is an extension to NTU 60 dataset. There are 114,480 videos collected with 120 action categories, performed by 106 subjects. Meanwhile, 32 collection setups with respect to the location and background are used to build the dataset. Two recommended protocols are presented: a) Cross-Subject (xsub): the data for training are collected from 53 subjects, while the testing data are from the other 53 subjects. b) Cross-Setup (xset): the training data uses even setup IDs, while testing data are odd setup IDs.
3) PKU Multi-Modality Dataset (PKUMMD) [28] is another large-scale benchmark with available skeleton data. Two subsets, Part I and Part II, are provided. PKUMMD Part I contains 1,076 long video sequences, with 20 action labels per video on average, and $\sim$20,000 instances are included in 51 action categories after trimming. Part II contains 2000 short video sequences with approximately seven instances each, focusing on the short-margin action detection task. It is more challenging due to the data noise and view variation.
4) Northwestern-UCLA (NW-UCLA) [84] contains 1,494 action samples, captured by three Kinect v1 cameras. Samples of 10 action categories are included, performed by 10 subjects. Following the recommended protocol, the training data are from the first two cameras while the testing data are from the other one.
4.2 Model Backbones
Different backbones are studied in previous skeleton SSL works, i.e., RNN-based, GCN-based, and Transformer-based models. RNNs treat the skeleton sequences as the temporal series and model temporal dependencies. However, it ignores the spatial structures of skeleton. Inspired by the natural topology structure of the human body, GCNs are widely explored to model spatial-temporal relationships. Recently, Transformer has been utilized to capture the long-temporal dependencies and has demonstrated remarkable results, owing to the attention mechanism.
Besides, some SSL works turn to other model backbones. Convolutional neural networks (CNNs) are utilized to process the skeleton sequence as a pseudo-2D image. Meanwhile, skeletons can also be treated as point clouds, and hence the Dynamic Graph CNN (DGCNN) [95] is also explored as the feature extractor.
4.3 Downstream Tasks
1) Action Recognition is the most common evaluation task for skeleton-based SSL works. Typically, two evaluation protocols are widely adopted. The first is the linear evaluation protocol, where a linear layer is added with the pre-trained model fixed. The other is fine-tuning protocol where the whole model is trained including the subsequent linear layer. Top-k accuracy metric is adopted in this task.
2) Action Retrieval aims to find the skeleton sequences that are similar or near-duplicates of a given query sequence. The metric precision, which is the proportion of retrieved relevant skeleton sequences in all retrieved entries, is reported.
3) Action Detection is a task that detects the start and end time of actions in an untrimmed skeleton sequence as well as its corresponding action label. It can also be referred to as the Temporal Action Localization or Action Segmentation. Following the previous works [96, 28], the Average Precision (mAP) at different temporal Intersection over Union (tIoU) thresholds between the predicted and the ground truth intervals is utilized as the metric.
4) Occluded Action Recognition focuses on the action understanding with occlusion, which is prevalent in human activity. Typically, the skeleton joints with low confidence or known to be occluded are set to zeros. The model is constrained to predict the correct label from corrupted skeleton sequences with top-k accuracy as the metric.
Figure 5: Action recognition performance of model over time for different SSL methodologies and different backbones. Left: NTU 60, Right: NTU 120, using cross-subject linear evaluation protocol.
4.4 Implementations Details
For the existing methods, we give priority to the reported results in their original paper. While for the reproduced results, we strictly keep the fairness of the comparison and implementation. Detailed settings of the reproduction experiments can be found in the supplementary material.
For our proposed method PCM3++, we adopt the three-layer Bi-GRU with the hidden dimension of $d=1024$ as the encoder backbone and follow the experiment settings including the data pre-processing and pre-training of our baseline method [71], which are consistent with other GRU-based works [10, 62] for comparison fairness. During the pre-training stage, the model is trained for 450 epochs with additional 150 epochs for post refinement. The batch size is 128. We set learning rate as 0.02, which is reduced to 0.002 at the 350th epoch. The SGD optimizer is adopted with a momentum of 0.9. $\tau$, $\tau_{q}$, and $\tau_{k}$ are set to 0.07, 0.1, and 0.05, respectively. $\lambda_{m}$ and $\lambda_{kd}$ are 40.0 and 1.0. We conduct all experiments on an NVIDIA RTX 4090 GPU.
4.5 Benchmark on Action Recognition Task
We benchmark different SSL methods on NTU-60 and NTU-120 benchmarks using the widely adopted action recognition task in Table LABEL:tab:main_comp. More comprehensive comparison on additional datasets can be found in the supplementary material. The feature dimension of the last layer before the classifier is provided, which often correlates with the upper capacity of the model. For the evaluation of other tasks, the community still lacks unified benchmarks. To this end, we reproduce some popular works with the same settings as our work to obtain comparable results in Sec. 4.6.
4.5.1 Comparison across Different Pretext Tasks
As shown in Fig. 5, context-based and generative learning pretext tasks are mainly studied in the earlier works first. However, these methods usually fail to achieve a satisfactory performance under linear evaluation due to the modeling of too many low-level features and the lack of effective design. Recently, MAMP [48] adopts MSM with large mask ratios and motion-aware reconstruction targets, obtaining a highly discriminative semantic feature space for both linear evaluation and fully fine-tuning protocol, which means that Transformer-based generative learning is promising.
Contrastive learning has always been a popular topic triggered by the success in the image field, e.g., SimCLR and MoCo. The model can learn a separable high-level representation space, making it dominant in learning linear representations. These works start from the exploiting the baseline algorithms [61, 62], and significantly boost the performance by applying stronger augmentations [60, 12, 67], effective training strategies [70, 73], cross-modal knowledge [9, 10], achieving rapid improvement in recent three years.
Besides, combining different pre-training paradigms [71, 78] for skeleton-based representation learning has demonstrated promising results recently. This indicates that the representations learned by different pre-training paradigms can be complementary and beneficial.
4.5.2 Comparison across Different Model Backbones
Different model backbones often employ different pre-training tasks. For example, Transformer often consumes extensive computational resources, making it difficult for contrastive learning, which often requires encoding multiple data views. In contrast, it naturally fits the MSM schema which reduces computational overhead by masking tokens with large ratios. On the other hand, GRU and GCN models are relatively more efficient in training, often adopting the contrastive learning pretext task. It is also noted that the GRU models are often with larger feature dimensions than GCN models due to the smaller induced computational graph and less GPU memory occupancy.
Generally, different from supervised training, GRU models have achieved state-of-the-art performance under linear evaluation as shown in Fig. 5 in SSL, and the Transformer models are also promising recently. For fully fine-tuning protocol, Transformer models are more popular due to the strong representation power. Remarkably, with self-supervised pre-training, a vanilla Transformer [10] has surpassed the GCN/Transformer with complex designs in supervised training [97, 98], demonstrating the significant role for alleviating the over-fitting problem and generalization capacity improvement of SSL.
[…]
Method xview xsub
1% data 10% data 1% data 10% data
Semi-Supervised Methods
X-CAR [80] - 78.2 - 76.1
MAC-Learning [81] - 78.5 - 74.2
Self-Supervised Methods
LongT GAN [7] - - 35.2 62.0
MS2L [13] - - 33.1 65.1
ISC [62] 38.1 72.5 35.7 65.9
CMD [10] 53.0 80.2 50.6 75.4
HiCo [11] 54.8 78.3 54.4 73.0
PCM3 [71] 53.1 82.8 53.8 77.1
PCM3++ 57.5 84.5 57.1 79.4
[…]
4.6 Benchmark on Different Downstream Tasks
To give a thorough evaluation of generalization capacity of current SSL models and our method PCM3++, we conduct extensive experiments and perform benchmarking on five downstream tasks, including action recognition, action retrieval, occluded recognition, action detection, and few-shot learning. For comparison, we choose the popular SSL methods mostly based on GRU model, which achieves a strong performance currently, while also ensuring a fair comparison with our GRU-based approach.
1) Skeleton-based Action Recognition. The following two evaluation settings are adopted.
$\bullet$ Unsupervised Learning Setting follows the linear evaluation protocol, adding a trainable linear classifier after the fixed pre-trained encoder. As shown in Table LABEL:tab:main_comp, PCM3++ achieves better or competitive results compared with other state-of-the-art methods on different evaluation protocols. Meanwhile, the results on PKUMMD dataset can be found in supplementary material, which also indicates the strong generalization capacity on the noisy data.
$\bullet$ Semi-supervised Learning Setting fine-tunes the whole model with only a portion of labeled data. This reflects the effectiveness in terms of avoiding over-fitting problem of the pretrained model. In addition to fine-tuning the whole model, we find the linear evaluation protocol tends to yield a better performance, especially for the small ratio of labeled data. As shown in Table II, PCM3++ obtains better performance than both the semi- and self-supervised methods with different proportions of training data. In contrast, previous works based on reconstruction [7] or contrastive learning [62, 10] solely can not achieve satisfactory results.
TABLE III: The results of action retrieval with joint stream.
Method NTU 60 (%) NTU 120 (%)
xsub xview xsub xset
LongT GAN [7] 39.1 48.1 31.5 35.5
ISC [62] 62.5 82.6 50.6 52.3
HiCLR [12] 60.6 73.1 46.0 46.0
CRRL [15] 60.7 75.2 - -
CMD [10] 70.6 85.4 58.3 60.9
HaLP [68] 65.8 83.6 55.8 59.0
HiCo [11] 67.9 84.4 55.9 58.7
MAMP [48] 62.0 70.0 - -
UmURL [75] 71.3 88.3 58.5 60.9
PCM3 [71] 73.7 88.8 63.1 66.8
PCM3++ 75.4 89.4 64.5 67.1
TABLE IV: The results of occluded action recognition. $\Delta_{\downarrow}$ represents the average performance degradation compared to that without occlusion.
Method Occluded NTU 60
Spatial Occ. (%) Temporal Occ. (%)
xsub xview $\Delta_{\downarrow}$ xsub xview $\Delta_{\downarrow}$
MoCo-GRU [49] 64.8 72.6 12.4 68.8 74.8 9.3
ISC [62] 62.8 70.6 14.1 68.9 76.8 7.9
CRRL [15] 56.8 61.4 11.6 61.0 66.2 7.1
CMD [10] 67.1 72.7 13.3 72.7 79.5 7.1
HiCo [11] 66.4 72.1 15.4 71.2 76.7 10.6
PCM3 [71] 80.8 87.0 3.3 77.6 86.1 5.4
PCM3++ 81.8 88.0 3.0 79.9 85.8 5.0
2) Skeleton-based Action Retrieval. Following previous work [14], a K-nearest neighbors (KNN) classifier ($k$=1) is adopted to retrieve the nearest training sample for each testing data in the representation space. The precision is reported as accuracy in Table III. Our method achieves a competitive performance compared with other latest methods. This indicates the highly distinguishable representations learned by our method. Meanwhile, it is found that contrastive learning based methods usually perform better than reconstruction-based methods, e.g., LongT GAN [7] and MAMP [48].
[…]
Method mAP@tIoU (%)
0.1 0.3 0.5
Supervised Traininig 63.4 61.4 53.4
Unsupervised Methods
MoCo-GRU [49] 68.2 67.2 63.4
CRRL [15] 57.6 55.7 52.1
ISC [62] 64.6 62.9 58.7
CMD [10] 73.7 72.8 68.4
HiCo [11] 51.8 50.8 45.8
[…]
PCM3 [71] 73.3 72.8 68.2
PCM3++ 75.5 74.6 69.8
TABLE VI: Comparison of few shot learning for skeleton-based action recognition. ($k$, $n$) represents $k$-way $n$-shot task.
Method xset (%) xsub (%)
(5, 1) (5, 5) (5, 1) (5, 5)
Supervised Training 64.5 81.2 67.3 85.2
Unsupervised Methods
MoCo-GRU [49] 58.5 78.7 61.1 81.3
CRRL [15] 58.3 77.1 55.2 75.0
ISC [62] 62.3 81.3 64.4 83.6
CMD [10] 64.9 82.8 68.6 85.7
HiCo [11] 64.6 81.7 67.1 84.2
PCM3 [71] 65.2 82.9 68.8 86.8
PCM3++ 66.7 84.8 69.9 87.3
3) Action Recognition with Occlusion. We evaluate the transfer ability of representations learned from the clean dataset to the action recognition with occluded data. We adopt the linear evaluation protocol. Following [100], we construct a synthetic occluded dataset on both spatial and temporal dimension. For spatial dimension, different body parts are randomly masked. For temporal dimension, we set a random block of frames to zeros. The testing set is constructed by the same masks across different methods, with a masking ratio of [0.3, 0.7].
As presented in Table IV, our method can capture underlying structures in the distorted data owing to the masked contrastive learning, and is good at dealing with the spatial occlusion. Meanwhile, this capacity extends well for temporal occlusion, demonstrating a desirable generalization capacity. However, due to the lack of modeling learning of occluded data, other methods are with poor robustness for occlusion.
4) Skeleton-based Action Detection. We follow the previous works [96, 28], and evaluate the short-term motion modeling capacity by action detection task. We only train the attached linear classifier (fully-connected layer) after the pre-trained encoder to predict frame-level categories and formulate the final proposal. The encoder is pre-trained on NTU 60 dataset and then we transfer the learned representations to the untrimmed PKUMMD Part I dataset. The mean average precision (mAP) of different actions is adopted as metric with different temporal overlapping ratios (tIoU) in Table V. Note that we provide a fully fine-tuned model (train both the encoder and classifier) without pre-training as the baseline. First, it can be found that the SSL pre-training can significantly improve the performance compared with the baseline model. Besides, our method integrates the clip-level representation contrastive learning and achieves the best scores compared with latest CMD [10], HiCo [11], and LAC [99], which ignore the clip-level representation modeling.
TABLE VII: Ablation study on the clip-level contrastive learning, the prompted training, and the post-distillation.
Method Recognition (%) Detection (%)
xsub xview Part I
w/o clip contrastive learning 83.9 90.4 73.3
w clip contrastive learning 84.1 90.8 75.5
w/o post-distillation 84.1 90.8 75.5
w post-distillation 84.8 91.0 75.5
TABLE VIII: Ablation study on the representation learning at different levels. Multi-Task denotes their simple combination of contrastive learning and masked prediction. $J,C,S$ represents the features of joint, clip, and sequence levels.
Method Representation Recognition Retrieval Recognition w Occlusion Detection Few Shot Learning
Acc. (%) Acc. (%) Acc. (%) mAP@0.1 (%) Acc. (%)
Masked Prediction J 14.7 64.4 11.7 36.1 43.5
Contrastive Learning S 87.3 84.5 76.6 72.5 61.4
Multi-Task J + S 87.5 85.1 77.4 71.9 60.2
PCM3 J + S 90.4 88.8 87.0 73.3 65.2
PCM3++ J + C + S 91.0 89.4 88.0 75.5 66.7
5) Unsupervised Few-shot Learning. In this task, we evaluate the performance of SSL as few-shot learners follow the previous work [101]. Specifically, the model is first pre-trained on NTU 60 dataset. Then, a simple classifier, e.g., Support Vector Machine (SVM) in our implementation, is fitted on the output features by the pre-trained encoder of the support data set. Finally, the adapted classifier along with the encoder is utilized to infer the query samples. We select 20 new categories in NTU 120111Specifically, the classes (index from 0) 60, 61, 66, 69, 72, 78, 79, 80, 84, 90, 91, 95, 96, 98, 99, 100, 102, 106, 108, 111, 113 114, 115 are selected. which are not seen in pre-training, as the support and query sets following [102]. In Table VI, our method generalizes well on the unseen categories and surpasses the supervised baseline and other methods notably, although there is no training specifically for unseen categories.
6) Compared with the Baseline Method. We evaluate and present the comparison results of our baseline model, PCM3 [71], on various downstream tasks as shown in Table LABEL:tab:main_comp-VI. As our new improvement, we integrate an effective clip-level contrastive learning scheme and present a novel post-distillation training strategy. These methodological advancements consistently improve the model’s representation quality and robustness on various downstream tasks, especially the semi-supervised action recognition and action detection performance.
To sum up, self-supervised learning provides a strong solution to boost various downstream tasks for versatile skeleton-based action understanding. Generally, a higher recognition performance roughly indicates a better representation space for most downstream tasks, as found in Table III and Table VI, but not equal to greater robustness to occlusion. Meanwhile, action detection is another crucial task for action analysis, which mainly focus on the short-term motion modeling. We design a novel asymmetric clip-level contrastive learning pretext task and effectively boost the detection performance. Remarkably, our proposed method achieves promising results on these tasks compared with other methods that ignore the versatile representation learning of different granularity.
4.7 Analysis of PCM3++
4.7.1 Ablation Study of Relevant Components
1) Analysis of clip contrastive learning. We add the temporal clips as the asymmetric positive sample of the anchor sequence, which further improves the representation quality as shown in Table VII. Remarkably, the clip-level contrastive learning significantly boost the action detection performance by promoting short-term modeling capacity.
2) Analysis of the versatile representation learning. As shown in Table VIII, we analyze the performance under different downstream tasks when adopting single paradigm or combining them. First, as discussed before, only employing the masked skeleton modeling cannot obtain a distinguishable representation space for GRU, leading to poor results in linear discrimination task. On the other hand, contrastive learning can achieve decent results on different downstream tasks by model the high-level semantics. However, it only learns the sequence-level representations, leaving the generalization capacity under-explored. Therefore, to achieve more representative features, we can combine different paradigms. However, simply combining them only shows mediocre improvement because it neglects the connection between the two tasks. In contrast, PCM3 [71] utilizes the synergy between the two tasks, and significantly improve the performance and generalization capacity. In addition, we integrate the clip-level representations and achieve further improvement.
3) Effect of the post-distillation design. As shown in Table VII, the post-distillation training strategy can bring further improvement slightly for the recognition task. It alleviates the false negative problem by removing the one-hot pseudo label in the contrastive learning, achieving a more compact representation space.
TABLE IX: Action recognition results using different backbones with joint stream.
Method Backbone xsub (%) xview (%)
ActCLR [67] GCN 80.9 86.7
CSTCN [91] GRU 83.1 88.7
HiCLR [12] Transformer 76.6 80.8
PCM3++ GCN 80.6 85.1
PCM3++ GRU 84.8 91.0
PCM3++ Transformer 80.2 84.9
4.7.2 More Analysis Results
1) Results of different backbones. We give more results with different backbones in Table IX. Specifically, for GCN and Transformer, we adopt ST-GCN [4] and DSTA-Net [98] following previous works [67, 12]. As we can see, our method shows good generalization capacity across different model backbones, and performs better or on-par compared with the latest SSL methods. Additional improvement can be potentially achieved by further searching for hyper-parameters on different backbones, which is ,however, not the focus of this paper, and we leave it for future work. We finally adopt GRU model in implementation for the more friendly memory usage of GPU and the higher performance.
2) Complexity Analysis. We give an analysis of space and computational complexities of our method for pre-training in Table X. As we can see, compared with other GRU-based and Transformer-based methods, our method achieves a significant performance improvement with an acceptable cost of the complexity. For the space complexity, the main additional cost is the reconstruction decoder $dec(\cdot)$, which takes up about 4M of space. For the computational complexity, the encoding process of different positive samples contributes most of the computational overhead.
TABLE X: FLOPs and Params results of different models.
Models Params $\downarrow$ FLOPs $\downarrow$ Accuracy
GL-Transformer [41] 214M 59.4G 83.8%
ISC [62] 106M 13.7G 85.2%
CMD [10] 99M 17.3G 86.9%
PCM3 [71] 103M 15.0G 90.4%
PCM3++ 103M 17.9G 91.0%
5 Conclusions and Future Directions
This paper presents a comprehensive survey on self-supervised skeleton-based action representation learning, where different literature is organized following the taxonomy of context-based, generative learning, and contrastive learning approaches. Then, a detailed benchmark and insightful discussions are provided, including the datasets, model backbones and pretext tasks. After reviewing existing works, we technically propose a novel and effective framework, for versatile skeleton-based action representation learning which is less explored before as a challenging topic. Revisiting the combination of contrastive learning and masked skeleton modeling paradigms, our method can achieve representation learning of different granularity by fully utilizing the novel spatial-temporal motion patterns. Extensive experiments on five downstream tasks demonstrate our superior performance and generalization capacity.
For the future research, the following pending issues deserve more attention:
  • Long-Term Motion Understanding. Existing methods utilize the trimmed video clip as training data, i.e., a video sample only contains one motion. This limits the long-term temporal reasoning capacity, especially for long videos containing multiple actions, e.g., Long-Term Action Anticipation, which should be paid more efforts.
  • Multi-Modal Learning. The exploration of multi-modal pre-training including human skeleton data is still insufficient. For example, as a useful complement to the skeleton data, RGB data can provide additional background knowledge that can boost action representation learning. Meanwhile, although valuable efforts have been made on the skeleton-text alignment [103, 104], stronger, larger-scale pre-training models are still urgently needed.
  • Versatile Representation Learning. As discussed in previous sections, more downstream tasks should be involved to fully explore and exploit the generalization capacity of SSL models. In addition to combining different pre-training paradigms analogous to this work, new frameworks can also be explored, e.g., Diffusion model [105], which is promising to handle both the skeleton-based generative and discriminative downstream tasks.
  • Towards Skeleton in the Wild. Existing methods are mostly evaluated in simplified and controlled environments and can suffer from serious noise caused by occlusion and view variation when deployed on a more diverse outdoor scenario. As an effective technique to improve model generalization, SSL is promising to boost skeleton representation robustness in the wild.
In summary, many abundant practices have emerged in skeleton-based SSL literature, and more valuable endeavours are expected to improve current SSL works and explore new directions.

### v2

@@
-Self-Supervised Skeleton Action Representation Learning: A Benchmark and Beyond
+Self-Supervised Skeleton-Based Action Representation Learning: A Benchmark and Beyond
 Abstract
-Self-supervised learning (SSL), which aims to learn meaningful prior representations from unlabeled data, has been proven effective for label-efficient skeleton-based action understanding.
-Different from the image domain, skeleton data possesses sparser spatial structures and diverse representation forms, with the absence of background clues and the additional temporal dimension.
-This presents the new challenges for the pretext task design of spatial-temporal motion representation learning.
-Recently, many endeavors have been made for skeleton-based SSL and remarkable progress has been achieved.
+Self-supervised learning (SSL), which aims to learn meaningful prior representations from unlabeled data, has been proven effective for skeleton-based action understanding.
+Different from the image domain, skeleton data possesses sparser spatial structures and diverse representation forms, with the absence of background clues and the additional temporal dimension, presenting new challenges for spatial-temporal motion pretext task design.
+Recently, many endeavors have been made for skeleton-based SSL, achieving remarkable progress.
 However, a systematic and thorough review is still lacking.
-In this paper, we conduct, for the first time, a comprehensive survey on self-supervised skeleton-based action representation learning, where various literature is organized according to their pre-training pretext task methodologies.
+In this paper, we conduct, for the first time, a comprehensive survey on self-supervised skeleton-based action representation learning.
 Following the taxonomy of context-based, generative learning, and contrastive learning approaches, we make a thorough review and benchmark of existing works and shed light on the future possible directions.
-Our investigation demonstrates that most SSL works rely on the single paradigm, learning representations of a single level, and are evaluated on the action recognition task solely, which leaves the generalization power of skeleton SSL models under-explored.
-To this end, a novel and effective SSL method for skeleton is further proposed, which integrates multiple pretext tasks to jointly learn versatile representations of different granularity, substantially boosting the generalization capacity for different downstream tasks.
-Extensive experiments under three large-scale datasets demonstrate that the proposed method achieves the superior generalization performance on various downstream tasks, including recognition, retrieval, detection, and few-shot learning.
-Index Terms:
+Remarkably, our investigation demonstrates that most SSL works rely on the single paradigm, learning representations of a single level, and are evaluated on the action recognition task solely, which leaves the generalization power of skeleton SSL models under-explored.
+To this end, a novel and effective SSL method for skeleton is further proposed, which integrates versatile representation learning objectives of different granularity, substantially boosting the generalization capacity for multiple skeleton downstream tasks.
+Extensive experiments under three large-scale datasets demonstrate our method achieves superior generalization performance on various downstream tasks, including recognition, retrieval, detection, and few-shot learning.
+keywords:
 Self-supervised learning, skeleton-based action understanding, contrastive learning, masked skeleton modeling
@@
 Owing to these desirable advantages, skeleton has attracted much attention in human action analysis.
-In the early works, many endeavors have been put into the supervised skeleton-based human activity understanding, e.g., recognition and detection [4, 5, 6].
+In the early works, many endeavors have been put into the supervised skeleton-based human activity understanding, e.g., recognition and detection [86, 63, 11, 59].
 However, these supervised methods heavily rely on huge amounts of labeled data, which requires time-consuming and expensive data annotation work, limiting the wide applications in the real world.
@@
 Based on the taxonomy of context-based, generative learning, and contrastive learning, we give a detailed analysis of the pretext task design and highlight the special consideration for skeleton data along with the corresponding challenges.
-  • We present a comprehensive benchmark of existing skeleton SSL works as well as a summary of the popular datasets and downstream tasks for SSL evaluation.
-Meanwhile, we provide insightful analysis from the perspective of model backbones and pre-training paradigms, and discuss the future possible directions.
   • Motivated by the limitations revealed by our survey, we explore skeleton-based versatile action representation learning to fully mine the generalization power of SSL models.
 An effective SSL schema is proposed, which integrates contrastive learning and MSM to jointly model the representations of different granularity, remarkably benefiting different downstream tasks.
-  • We perform rigorous quantitative experiments to study the generalization efficacy of current skeleton-based SSL methods across five downstream tasks, including recognition, retrieval, detection, and few-shot learning, on both uncorrupted and corrupted skeletons.
-Under the proposed SSL method, the promising results are achieved for versatile action representation learning.
+  • We present a comprehensive benchmark of existing skeleton SSL works as well as a summary of the popular datasets and downstream tasks as well as the insightful analysis from the perspective of model backbones, pre-training paradigms, and future possible directions.
+Finally, we demonstrate promising performance of our proposed method on five downstream tasks for versatile action representation learning.
 The remaining sections are organized as follows: We first present a thorough review in Sec.
@@
 5.
-2 Review on Self-Supervised Skeleton-Based Action Representation Learning
+2 Review on Skeleton-Based Action Representation SSL
 Generally, a two-stage paradigm is utilized in skeleton-based SSL, i.e., pre-training on pretext tasks first and then fine-tuning on downstream tasks.
@@
 The survey is structured around three dimensions: skeleton data collection, SSL pretext design, and SSL downstream task evaluation, providing a comprehensive review.
-2.1 Skeleton Data as Human Representation
+2.1 Human Skeleton Representation
 Skeleton represents a human motion as a sequence of the body keypoint coordinates over time.
@@
 Therefore, learning view-invariant representations should be beneficial for action recognition task, and has been widely studied in the supervised skeleton-based action recognition task [19].
-Li et al[31] proposed a view classification pretext task for unsupervised action representation learning.
+Li et al.
+[37] proposed a view classification pretext task for unsupervised action representation learning.
 Specifically, as shown in Fig.
@@
 Resort to this, the encoder can learn the features insensitive to view in an adversarial manner.
-Likewise, Paoletti et al[33] adopted the rotation prediction, i.e., pitch angle, yaw angle, and roll angle, as well as a gradient reverse operation to achieve the viewpoint-invariance learning.
+Likewise, Paoletti et al.
+[55] adopted the rotation prediction, i.e., pitch angle, yaw angle, and roll angle, as well as a gradient reverse operation to achieve the viewpoint-invariance learning.
 There are also some subsequent SSL works [87, 53, 3] focusing on learning view-invariant representations.
@@
 3 (c).
-Cheng et al[40] designed a movement direction prediction task.
+Cheng et al.
+[12] designed a movement direction prediction task.
 The model is constrained to estimate the direction of the instantaneous joint velocity, i.e., whether the targeted joints are moving in a positive direction.
-Based on this, Kim et al[41] proposed a multi-interval pose displacement strategy to encourage both local and global attention learning.
+Based on this, Kim et al.
+[33] proposed a multi-interval pose displacement strategy to encourage both local and global attention learning.
 The key idea is to predict the motion direction and magnitude of the central joint and other joints during a random interval.
-On the other hand, Yang et al[42] introduced more motion prior information as the pseudo label, i.e., intra-joint motion variance, inter-joint motion covariance, intra-frame joint angle, and inter-frame motion deviation.
+On the other hand, Yang et al.
+[91] introduced more motion prior information as the pseudo label, i.e., intra-joint motion variance, inter-joint motion covariance, intra-frame joint angle, and inter-frame motion deviation.
 The model is pre-trained by a regression task to predict this prior knowledge.
@@
 Then, denoising AE (DAE) is studied for alleviating this shortcut [69, 54].
-For example, Nie et al[45] applied a series of view corruptions and constrained the model to reconstruct the clean data to disentangle the view and pose features.
+For example, Nie et al.
+[54] applied a series of view corruptions and constrained the model to reconstruct the clean data to disentangle the view and pose features.
 On the other hand, inspired by the remarkable success of masked language/image modeling (MLM, MIM), which can be regarded as a variant of DAE, researchers have aroused a surge of interest in exploring masked skeleton modeling (MSM) for human action representation learning.
@@
 Some works [78] propose to perform temporally reverse reconstruction to learn the skeleton dynamics instead of trivial representations that just remember the input.
-Instead of directly reconstructing the input, Mao et al[48] proposed to reconstruct the motion in MSM, i.e., the difference of the corresponding joints between adjacent frames.
+Instead of directly reconstructing the input, Mao et al.
+[51] proposed to reconstruct the motion in MSM, i.e., the difference of the corresponding joints between adjacent frames.
 The work [90] formulates the skeleton as an unordered 3D point cloud and maps the 3D data onto color space.
@@
 Therefore, the model can learn the spatial relation and temporal dependency by reconstructing the color of the skeleton cloud.
-2.2.3 Contrastive-Learning-Based Methods
+2.2.3 Contrastive-Learning Methods
 Contrastive learning has been proven effective for different data modalities, e.g., images, point cloud, as well as skeleton data.
@@
 Specifically, different augmentations are applied to the skeleton $x$ to generate the positive pair $(x_{q},x_{k})$, while the negative sample features $m_{i}$ are stored in a memory queue.
-The model is constrained to retrieve the positives among the negative samples, optimizing the following InfoNCE objective:
-$\mathcal{L}_{Info}(z_{q},z_{k})=-\log\frac{\exp(z_{q}\cdot z_{k}/\tau)}{\exp(z_{q}\cdot z_{k}/\tau)+\sum_{i=1}{\exp(z_{q}\cdot m_{i}/\tau)}},$ (2)
+The model is constrained to retrieve the positives among the negative samples, optimizing the following InfoNCE objective $\mathcal{L}_{Info}(z_{q},z_{k})$:
+$\mathcal{L}=-\log\frac{\exp(z_{q}\cdot z_{k}/\tau)}{\exp(z_{q}\cdot z_{k}/\tau)+\sum_{i=1}{\exp(z_{q}\cdot m_{i}/\tau)}},$ (2)
 where the $z_{q}$/$z_{k}$ is the query/key embedding encoded by the query/key model and $\tau$ is the temperature hyper-parameter.
@@
 Therefore, the earlier works mainly focus on exploring practical spatial-temporal augmentations for skeleton contrastive learning [60, 72, 20].
-For example, Rao et al[61] proposed a series of augmentations including Rotation, Shear, Reverse, Gaussian Noise, Gaussian Blur, Joint Mask and Channel Mask, some of which are used as the default basic augmentations in the future research.
+For example, Rao et al.
+[60] proposed a series of augmentations including Rotation, Shear, Reverse, Gaussian Noise, Gaussian Blur, Joint Mask and Channel Mask, some of which are used as the default basic augmentations in the future research.
 Notably, the optimal augmentations for different backbones are usually different.
@@
 Generally, this difference arises on account of different modeling mechanisms and model capacities, which further increases the difficulty of the augmentation design and selection.
-Inspired by the success of mixing-based augmentations for images, Chen et al[64] proposed SkeleMixCLR equipped with SkeleMix augmentation which combines the topological information of different skeleton sequences.
+Inspired by the success of mixing-based augmentations for images, Chen et al.
+[94] proposed SkeleMixCLR equipped with SkeleMix augmentation which combines the topological information of different skeleton sequences.
 To learn from this mixed skeleton, SkeleMixCLR obtains the corresponding part level and the whole-body level features and pursues the local-global invariance.
@@
 Instead of the one-hot target, this objective utilizes the similarity distribution as the soft target to guide consistency learning, which can be viewed as a self-distillation process.
-Further, Zhang et al[12] introduced more strong augmentations, e.g., randomly dropping the skeleton edges and joints, and proposed a hierarchical contrastive learning framework.
+Further, Zhang et al.
+[97] introduced more strong augmentations, e.g., randomly dropping the skeleton edges and joints, and proposed a hierarchical contrastive learning framework.
 It performs a decoupled progressive augmentation invariance learning by optimizing the consistency only between the augmented samples with adjacent strength, where the weakly augmented branch serves as the mimic target of the strongly augmented branch.
 These works show that the weakly augmented views can effectively guide the learning of the corresponding strongly augmented samples, leading to more stable representation learning and improvement.
-In addition, Lin et al[67] explicitly distinguished the static regions and motion regions, namely, actionlet, in human skeleton, and introduced a motion-aware augmentation strategy.
+In addition, Lin et al.
+[41] explicitly distinguished the static regions and motion regions, namely, actionlet, in human skeleton, and introduced a motion-aware augmentation strategy.
 By mining the actionlet in an unsupervised manner, the semantic-reserving augmentations are employed for actionlet regions, while the noise perturbations for non-actionlet regions, avoiding the over-distortion problem.
@@
 Overall, although downstream tasks are often unknowable in pre-training, these methods can obtain a potentially beneficial representation space by introducing different prior knowledge.
-3 The Proposed Method: Prompted Contrast with Masked Motion Modeling
+3 The Proposed Method
 3.1 Motivation
@@
 Meanwhile, the evaluation of proposed PCM3++ is reported.
-TABLE I: Summary of popular skeleton datasets. Marker denotes the dataset is collected by Marker-MoCap methods.
-Dataset’Year Instance Class. Sub. View Joint Marker
-CMU [82]’03 2235 45 144 - 31
-SBU [83]’12 300 8 7 - 15
-NW-UCLA [84]’14 1,494 10 10 3 20
-UWA3D [85]’15 1,075 30 10 5 15
-SYSU [86]’15 480 12 40 1 20
-NTU 60 [26]’16 56,880 60 40 80 25
-PKUMMD [28]’17 21,545 51 66 3 25
-Kinetics [87]’17 $\sim$260k 400 - - 18
-TSU [88]’19 16,115 31 18 - 17
-NTU 120 [27]’19 114,480 120 106 155 25
-UAV [89]’21 22,476 155 119 17
-BABEL-60 [90]’21 45473 60 - - 25
-BABEL-120 [90]’21 48,978 120 - - 25
+Table 1: Summary of popular skeleton datasets. * denotes the collected by Marker-MoCap methods.
+Dataset Instance Class Sub. View Joint
+CMU* [13] 2235 45 144 - 31
+SBU [92] 300 8 7 - 15
+NW-UCLA [76] 1,494 10 10 3 20
+UWA3D [58] 1,075 30 10 5 15
+SYSU [28] 480 12 40 1 20
+NTU 60 [62] 56,880 60 40 80 25
+PKUMMD [43] 21,545 51 66 3 25
+Kinetics [32] $\sim$260k 400 - - 18
+TSU [14] 16,115 31 18 - 17
+NTU 120 [44] 114,480 120 106 155 25
+UAV [39] 22,476 155 119 17
+BABEL* [56] 48,978 120 - - 25
 4.1 Datasets
-We first summarize the popular datasets for skeleton action understanding in Table I. Among them, the following datasets are widely used in skeleton-based SSL evaluation.
+We first summarize the popular datasets for skeleton action understanding in Table 1. Among them, the following datasets are widely used in skeleton-based SSL evaluation.
 1) NTU RGB+D 60 Dataset (NTU 60) [62] is the most popular skeleton dataset.
@@
 It is more challenging due to the data noise and view variation.
-4) Northwestern-UCLA (NW-UCLA) [84] contains 1,494 action samples, captured by three Kinect v1 cameras.
-Samples of 10 action categories are included, performed by 10 subjects.
-Following the recommended protocol, the training data are from the first two cameras while the testing data are from the other one.
 4.2 Model Backbones
@@
 4.3 Downstream Tasks
-1) Action Recognition is the most common evaluation task for skeleton-based SSL works.
+1) Action Recognition is the most common SSL evaluation task.
 Typically, two evaluation protocols are widely adopted.
@@
 4.4 Implementations Details
+For the proposed method, we follow the implementation and experiment settings of previous works [98, 52].
 For the existing methods, we give priority to the reported results in their original paper.
-While for the reproduced results, we strictly keep the fairness of the comparison and implementation.
-Detailed settings of the reproduction experiments can be found in the supplementary material.
-For our proposed method PCM3++, we adopt the three-layer Bi-GRU with the hidden dimension of $d=1024$ as the encoder backbone and follow the experiment settings including the data pre-processing and pre-training of our baseline method [71], which are consistent with other GRU-based works [10, 62] for comparison fairness.
-During the pre-training stage, the model is trained for 450 epochs with additional 150 epochs for post refinement.
-The batch size is 128.
-We set learning rate as 0.02, which is reduced to 0.002 at the 350th epoch.
-The SGD optimizer is adopted with a momentum of 0.9.
-$\tau$, $\tau_{q}$, and $\tau_{k}$ are set to 0.07, 0.1, and 0.05, respectively.
-$\lambda_{m}$ and $\lambda_{kd}$ are 40.0 and 1.0.
-We conduct all experiments on an NVIDIA RTX 4090 GPU.
+Meanwhile, we reproduce some results to keep the fairness of the comparison and implementation.
+Details can be found in Supp.
 4.5 Benchmark on Action Recognition Task
@@
 4.6.
-4.5.1 Comparison across Different Pretext Tasks
+1) Comparison across Different Pretext Tasks.
 As shown in Fig.
@@
 However, these methods usually fail to achieve a satisfactory performance under linear evaluation due to the modeling of too many low-level features and the lack of effective design.
-Recently, MAMP [48] adopts MSM with large mask ratios and motion-aware reconstruction targets, obtaining a highly discriminative semantic feature space for both linear evaluation and fully fine-tuning protocol, which means that Transformer-based generative learning is promising.
+Recently, MAMP [51] adopts MSM with large mask ratios and motion-aware reconstruction targets, obtaining a highly discriminative semantic feature space for both linear evaluation and fully fine-tuning protocol, demonstrating promising results for Transformer-based generative learning.
 Contrastive learning has always been a popular topic triggered by the success in the image field, e.g., SimCLR and MoCo.
@@
 This indicates that the representations learned by different pre-training paradigms can be complementary and beneficial.
-4.5.2 Comparison across Different Model Backbones
+2) Comparison across Different Model Backbones.
 Different model backbones often employ different pre-training tasks.
@@
 Method xview xsub
-1% data 10% data 1% data 10% data
+1% 10% 1% 10%
 Semi-Supervised Methods
@@
 1) Skeleton-based Action Recognition. The following two evaluation settings are adopted.
-$\bullet$ Unsupervised Learning Setting follows the linear evaluation protocol, adding a trainable linear classifier after the fixed pre-trained encoder.
-As shown in Table LABEL:tab:main_comp, PCM3++ achieves better or competitive results compared with other state-of-the-art methods on different evaluation protocols.
+Firstly, the unsupervised learning results using linear evaluation protocol is shown in Table LABEL:tab:main_comp, and PCM3++ achieves better or competitive results compared with other state-of-the-art methods on different evaluation protocols.
 Meanwhile, the results on PKUMMD dataset can be found in supplementary material, which also indicates the strong generalization capacity on the noisy data.
-$\bullet$ Semi-supervised Learning Setting fine-tunes the whole model with only a portion of labeled data.
-This reflects the effectiveness in terms of avoiding over-fitting problem of the pretrained model.
-In addition to fine-tuning the whole model, we find the linear evaluation protocol tends to yield a better performance, especially for the small ratio of labeled data.
-As shown in Table II, PCM3++ obtains better performance than both the semi- and self-supervised methods with different proportions of training data.
+In addition, we also report the semi-supervised setting results, which fine-tunes the whole model/classifier with only a portion of labeled data.
+As shown in Table 2, PCM3++ obtains better performance than both the semi- and self-supervised methods with different proportions of training data, and only fine-tuning the classifier tends to be better with extremely little data.
 In contrast, previous works based on reconstruction [100] or contrastive learning [72, 52] solely can not achieve satisfactory results.
-TABLE III: The results of action retrieval with joint stream.
+Table 3: The results of action retrieval with joint stream.
 Method NTU 60 (%) NTU 120 (%)
 xsub xview xsub xset
-LongT GAN [7] 39.1 48.1 31.5 35.5
 ISC [72] 62.5 82.6 50.6 52.3
@@
 PCM3++ 75.4 89.4 64.5 67.1
-TABLE IV: The results of occluded action recognition. $\Delta_{\downarrow}$ represents the average performance degradation compared to that without occlusion.
+Table 4: The results of occluded action recognition.
 Method Occluded NTU 60
-Spatial Occ. (%) Temporal Occ. (%)
-xsub xview $\Delta_{\downarrow}$ xsub xview $\Delta_{\downarrow}$
-MoCo-GRU [49] 64.8 72.6 12.4 68.8 74.8 9.3
-ISC [62] 62.8 70.6 14.1 68.9 76.8 7.9
-CRRL [15] 56.8 61.4 11.6 61.0 66.2 7.1
-CMD [10] 67.1 72.7 13.3 72.7 79.5 7.1
-HiCo [11] 66.4 72.1 15.4 71.2 76.7 10.6
-PCM3 [71] 80.8 87.0 3.3 77.6 86.1 5.4
-PCM3++ 81.8 88.0 3.0 79.9 85.8 5.0
+Spatial Occ. Temporal Occ.
+xsub xview xsub xview
+MoCo-GRU [26] 64.8 72.6 68.8 74.8
+ISC [72] 62.8 70.6 68.9 76.8
+CRRL [78] 56.8 61.4 61.0 66.2
+CMD [52] 67.1 72.7 72.7 79.5
+HiCo [16] 66.4 72.1 71.2 76.7
+PCM3 [98] 80.8 87.0 77.6 86.1
+PCM3++ 81.8 88.0 79.9 85.8
 2) Skeleton-based Action Retrieval.
 Following previous work [68], a K-nearest neighbors (KNN) classifier ($k$=1) is adopted to retrieve the nearest training sample for each testing data in the representation space.
-The precision is reported as accuracy in Table III.
+The precision is reported as accuracy in Table 3.
 Our method achieves a competitive performance compared with other latest methods.
@@
 Meanwhile, it is found that contrastive learning based methods usually perform better than reconstruction-based methods, e.g., LongT GAN [100] and MAMP [51].
-[…]
+Table 5: Action detection results on PKUMMD Part I benchmark under linear evaluation protocol.
 Method mAP@tIoU (%)
 0.1 0.3 0.5
-Supervised Traininig 63.4 61.4 53.4
+Supervise 63.4 61.4 53.4
 Unsupervised Methods
@@
 HiCo [16] 51.8 50.8 45.8
-[…]
+LAC [88] 55.2 - -
 PCM3 [98] 73.3 72.8 68.2
 PCM3++ 75.5 74.6 69.8
-TABLE VI: Comparison of few shot learning for skeleton-based action recognition. ($k$, $n$) represents $k$-way $n$-shot task.
+Table 6: Comparison of few shot skeleton-based action recognition. ($k$, $n$) represents $k$-way $n$-shot.
 Method xset (%) xsub (%)
 (5, 1) (5, 5) (5, 1) (5, 5)
-Supervised Training 64.5 81.2 67.3 85.2
+Supervise 64.5 81.2 67.3 85.2
 Unsupervised Methods
@@
 The testing set is constructed by the same masks across different methods, with a masking ratio of [0.3, 0.7].
-As presented in Table IV, our method can capture underlying structures in the distorted data owing to the masked contrastive learning, and is good at dealing with the spatial occlusion.
+As presented in Table 4, our method can capture underlying structures in the distorted data owing to the masked contrastive learning, and is good at dealing with the spatial occlusion.
 Meanwhile, this capacity extends well for temporal occlusion, demonstrating a desirable generalization capacity.
@@
 We follow the previous works [49, 43], and evaluate the short-term motion modeling capacity by action detection task.
-We only train the attached linear classifier (fully-connected layer) after the pre-trained encoder to predict frame-level categories and formulate the final proposal.
-The encoder is pre-trained on NTU 60 dataset and then we transfer the learned representations to the untrimmed PKUMMD Part I dataset.
-The mean average precision (mAP) of different actions is adopted as metric with different temporal overlapping ratios (tIoU) in Table V.
-Note that we provide a fully fine-tuned model (train both the encoder and classifier) without pre-training as the baseline.
-First, it can be found that the SSL pre-training can significantly improve the performance compared with the baseline model.
+We train the attached linear classifier to predict frame-level categories to produce the final proposal.
+The encoder is pre-trained on NTU 60 dataset and then we fix it, transferring the learned representations to untrimmed PKUMMD Part I dataset.
+The mean average precision (mAP) of different actions is adopted as metric with different temporal overlapping ratios (tIoU) in Table 5.
+We provide a fully fine-tuned model (train both the encoder and classifier) without pre-training as the baseline.
+First, it can be found the SSL pre-training can significantly improve the performance compared with the baseline model.
 Besides, our method integrates the clip-level representation contrastive learning and achieves the best scores compared with latest CMD [52], HiCo [16], and LAC [88], which ignore the clip-level representation modeling.
-TABLE VII: Ablation study on the clip-level contrastive learning, the prompted training, and the post-distillation.
-Method Recognition (%) Detection (%)
+Table 7: Ablation study on the clip-level contrastive learning and the post-distillation.
+Method Recognition Detection
 xsub xview Part I
@@
 w post-distillation 84.8 91.0 75.5
-TABLE VIII: Ablation study on the representation learning at different levels.
-Multi-Task denotes their simple combination of contrastive learning and masked prediction.
-$J,C,S$ represents the features of joint, clip, and sequence levels.
-Method Representation Recognition Retrieval Recognition w Occlusion Detection Few Shot Learning
-Acc. (%) Acc. (%) Acc. (%) mAP@0.1 (%) Acc. (%)
-Masked Prediction J 14.7 64.4 11.7 36.1 43.5
-Contrastive Learning S 87.3 84.5 76.6 72.5 61.4
-Multi-Task J + S 87.5 85.1 77.4 71.9 60.2
-PCM3 J + S 90.4 88.8 87.0 73.3 65.2
-PCM3++ J + C + S 91.0 89.4 88.0 75.5 66.7
 5) Unsupervised Few-shot Learning.
@@
 which are not seen in pre-training, as the support and query sets following [46].
-In Table VI, our method generalizes well on the unseen categories and surpasses the supervised baseline and other methods notably, although there is no training specifically for unseen categories.
+In Table 6, our method generalizes well on the unseen categories and surpasses the supervised baseline and other methods notably, although there is no training for unseen categories.
 6) Compared with the Baseline Method.
-We evaluate and present the comparison results of our baseline model, PCM3 [71], on various downstream tasks as shown in Table LABEL:tab:main_comp-VI.
+We evaluate and present the comparison results of our baseline model, PCM3 [98], on various downstream tasks as shown in Table LABEL:tab:main_comp-6.
 As our new improvement, we integrate an effective clip-level contrastive learning scheme and present a novel post-distillation training strategy.
@@
 To sum up, self-supervised learning provides a strong solution to boost various downstream tasks for versatile skeleton-based action understanding.
-Generally, a higher recognition performance roughly indicates a better representation space for most downstream tasks, as found in Table III and Table VI, but not equal to greater robustness to occlusion.
+Generally, a higher recognition performance roughly indicates a better representation space for most downstream tasks, as found in Table 3 and Table 6, but not equal to greater robustness to occlusion.
 Meanwhile, action detection is another crucial task for action analysis, which mainly focus on the short-term motion modeling.
@@
 4.7 Analysis of PCM3++
-4.7.1 Ablation Study of Relevant Components
 1) Analysis of clip contrastive learning.
-We add the temporal clips as the asymmetric positive sample of the anchor sequence, which further improves the representation quality as shown in Table VII.
+We add the temporal clips as the asymmetric positive sample of the anchor sequence, which further improves the representation quality as shown in Table 7.
 Remarkably, the clip-level contrastive learning significantly boost the action detection performance by promoting short-term modeling capacity.
-2) Analysis of the versatile representation learning.
-As shown in Table VIII, we analyze the performance under different downstream tasks when adopting single paradigm or combining them.
-First, as discussed before, only employing the masked skeleton modeling cannot obtain a distinguishable representation space for GRU, leading to poor results in linear discrimination task.
-On the other hand, contrastive learning can achieve decent results on different downstream tasks by model the high-level semantics.
-However, it only learns the sequence-level representations, leaving the generalization capacity under-explored.
-Therefore, to achieve more representative features, we can combine different paradigms.
-However, simply combining them only shows mediocre improvement because it neglects the connection between the two tasks.
-In contrast, PCM3 [71] utilizes the synergy between the two tasks, and significantly improve the performance and generalization capacity.
-In addition, we integrate the clip-level representations and achieve further improvement.
-3) Effect of the post-distillation design.
-As shown in Table VII, the post-distillation training strategy can bring further improvement slightly for the recognition task.
+Our designs boost versatile representation learning of different granularity and the effect on more downstream tasks can be found in Supp.
+2) Effect of the post-distillation design.
+As shown in Table 7, the post-distillation training strategy can bring further improvement slightly for the recognition task.
 It alleviates the false negative problem by removing the one-hot pseudo label in the contrastive learning, achieving a more compact representation space.
-TABLE IX: Action recognition results using different backbones with joint stream.
-Method Backbone xsub (%) xview (%)
+Table 8: Results of different backbones.
+Method Backbone xsub xview
 ActCLR [41] GCN 80.9 86.7
@@
 PCM3++ Transformer 80.2 84.9
-4.7.2 More Analysis Results
-1) Results of different backbones.
-We give more results with different backbones in Table IX.
-Specifically, for GCN and Transformer, we adopt ST-GCN [4] and DSTA-Net [98] following previous works [67, 12].
-As we can see, our method shows good generalization capacity across different model backbones, and performs better or on-par compared with the latest SSL methods.
-Additional improvement can be potentially achieved by further searching for hyper-parameters on different backbones, which is ,however, not the focus of this paper, and we leave it for future work.
-We finally adopt GRU model in implementation for the more friendly memory usage of GPU and the higher performance.
-2) Complexity Analysis.
-We give an analysis of space and computational complexities of our method for pre-training in Table X.
-As we can see, compared with other GRU-based and Transformer-based methods, our method achieves a significant performance improvement with an acceptable cost of the complexity.
-For the space complexity, the main additional cost is the reconstruction decoder $dec(\cdot)$, which takes up about 4M of space.
-For the computational complexity, the encoding process of different positive samples contributes most of the computational overhead.
-TABLE X: FLOPs and Params results of different models.
-Models Params $\downarrow$ FLOPs $\downarrow$ Accuracy
+Table 9: FLOPs and Params results.
+[…]
 GL-Transformer [33] 214M 59.4G 83.8%
@@
 PCM3++ 103M 17.9G 91.0%
+3) Results of different backbones.
+We give more results with different backbones in Table 8.
+Specifically, for GCN and Transformer, we adopt ST-GCN [86] and DSTA-Net [64] following previous works [41, 97].
+As we can see, our method shows good generalization capacity across different model backbones, and performs better or on-par compared with the latest SSL methods.
+Additional improvement can be potentially achieved by further searching for hyper-parameters on different backbones, which is, however, not the focus of this paper, and we leave it for future work.
+We finally adopt GRU model in implementation for the more friendly memory usage of GPU and the higher performance.
+4) Complexity Analysis.
+We give an analysis of space and computational complexities of our method for pre-training in Table 9.
+Compared with other GRU-based and Transformer-based methods, our method achieves a significant performance improvement with an acceptable cost of the complexity.
+For the space complexity, the main extra cost is the reconstruction decoder (4M of space).
+For the computational complexity, the encoding process of different positive samples contributes most of the overhead.
 5 Conclusions and Future Directions
-This paper presents a comprehensive survey on self-supervised skeleton-based action representation learning, where different literature is organized following the taxonomy of context-based, generative learning, and contrastive learning approaches.
-Then, a detailed benchmark and insightful discussions are provided, including the datasets, model backbones and pretext tasks.
+This paper presents a comprehensive survey on skeleton-based action representation SSL, where different literature is organized following the taxonomy of context-based, generative learning, and contrastive learning approaches.
 After reviewing existing works, we technically propose a novel and effective framework, for versatile skeleton-based action representation learning which is less explored before as a challenging topic.
-Revisiting the combination of contrastive learning and masked skeleton modeling paradigms, our method can achieve representation learning of different granularity by fully utilizing the novel spatial-temporal motion patterns.
-Extensive experiments on five downstream tasks demonstrate our superior performance and generalization capacity.
+Finally, a detailed benchmark and insightful discussions are provided, including datasets, model backbones and pretext tasks.
+Meanwhile, we demonstrate the superior generalization performance of our method under different downstream tasks.
 For the future research, the following pending issues deserve more attention:
   • Long-Term Motion Understanding.
-Existing methods utilize the trimmed video clip as training data, i.e., a video sample only contains one motion.
-This limits the long-term temporal reasoning capacity, especially for long videos containing multiple actions, e.g., Long-Term Action Anticipation, which should be paid more efforts.
+Existing methods utilize the trimmed video clip, only containing one motion, as training data.
+This limits the long-term temporal reasoning capacity containing multiple actions, e.g., Long-Term Action Anticipation Task.
   • Multi-Modal Learning.
-The exploration of multi-modal pre-training including human skeleton data is still insufficient.
-For example, as a useful complement to the skeleton data, RGB data can provide additional background knowledge that can boost action representation learning.
-Meanwhile, although valuable efforts have been made on the skeleton-text alignment [103, 104], stronger, larger-scale pre-training models are still urgently needed.
+The exploration of multi-modal pre-training including human skeleton data is still insufficient, e.g., text, infrared data, and RGB images.
+For example, as a useful complement to skeleton, RGB data can provide additional background knowledge that can boost action representation learning.
+  • Towards Skeleton in the Wild.
+Existing methods are evaluated in simplified and controlled environments and can suffer from serious noise, e.g., occlusion and view variation when deployed on the outdoor scenario.
+On this point, SSL is promising to boost skeleton representation robustness in the wild.
   • Versatile Representation Learning.
-As discussed in previous sections, more downstream tasks should be involved to fully explore and exploit the generalization capacity of SSL models.
-In addition to combining different pre-training paradigms analogous to this work, new frameworks can also be explored, e.g., Diffusion model [105], which is promising to handle both the skeleton-based generative and discriminative downstream tasks.
-  • Towards Skeleton in the Wild.
-Existing methods are mostly evaluated in simplified and controlled environments and can suffer from serious noise caused by occlusion and view variation when deployed on a more diverse outdoor scenario.
-As an effective technique to improve model generalization, SSL is promising to boost skeleton representation robustness in the wild.
-In summary, many abundant practices have emerged in skeleton-based SSL literature, and more valuable endeavours are expected to improve current SSL works and explore new directions.
+As discussed in previous sections, most works only focus on the recognition task, leaving the generation capacity across different tasks of SSL models under-explored.
+Meanwhile, new frameworks can also be explored, e.g., Diffusion model [27], which is promising to handle both generative and discriminative downstream tasks.
+Declarations
+  • Competing interests.
+This work was supported in part by the National Natural Science Foundation of China under Grant No.62172020, and in part by the Key Laboratory of Science, Technology and Standard in Press Industry (Key Laboratory of Intelligent Press Media Technology).
+  • Data availability.
+The adopted NTU-RGB+D [62, 44] and PKU-MMD [43] dataset are popular public benchmarks in skeleton-based action understanding.
+The code for data processing can be found in https://github.com/JHang2020/PCM3.

### v3

@@
 1 Introduction
-Human activity understanding is an essential topic in the research of computer vision due to its wide applications in real life, such as human-robotics interaction [36], autonomous driving [5], and healthcare [47].
+Human activity understanding is an essential topic in the research of computer vision due to its wide applications in real life, such as human-robotics interaction [lee2020real], autonomous driving [camara2020pedestrian], and healthcare [lopez2019human].
 Among the different data modalities for actions, skeletons represent the human body by 3D coordinates of key body joints, which are lightweight, compact, and more robust to changes of view and background.
 Owing to these desirable advantages, skeleton has attracted much attention in human action analysis.
-In the early works, many endeavors have been put into the supervised skeleton-based human activity understanding, e.g., recognition and detection [86, 63, 11, 59].
+In the early works, many endeavors have been put into the supervised skeleton-based human activity understanding, e.g., recognition and detection [yan2018spatial, shi2019two, cheng2020skeleton, rao2024hierarchical].
 However, these supervised methods heavily rely on huge amounts of labeled data, which requires time-consuming and expensive data annotation work, limiting the wide applications in the real world.
@@
 We believe our extensive work can benefit the research community and bring rich insights for future work.
-Based on our literature review, it is noticed that most SSL methods for skeleton focus on the single paradigm, learning representations of single granularity, e.g., joint-level features (by masked skeleton modeling, MSM) [100, 80] or sequence-level features (by contrastive learning) [38, 52, 16, 97].
+Based on our literature review, it is noticed that most SSL methods for skeleton focus on the single paradigm, learning representations of single granularity, e.g., joint-level features (by masked skeleton modeling, MSM) [zheng2018unsupervised, wu2022skeletonmae] or sequence-level features (by contrastive learning) [li20213d, mao2022cmd, hico2023, zhang2022hiclr].
 This limits the generalization capacity of the model to more downstream tasks.
-Although some works [40, 68, 78] make valuable efforts to combine different paradigms, they only achieve mediocre improvement due to the inherent gap of feature modeling mechanisms between the contrastive learning and masked modeling [57, 23].
+Although some works [lin2020ms2l, su2020predict, wang2022contrast] make valuable efforts to combine different paradigms, they only achieve mediocre improvement due to the inherent gap of feature modeling mechanisms between the contrastive learning and masked modeling [qi2023contrast, gui2023survey].
 To this end, a novel SSL approach for skeleton is proposed to fully boost the generalization capacity of SSL model, which integrates the contrastive learning and MSM to learn the joint, clip, and sequence level representations jointly.
 Specifically, we fully utilize the novel motion pattern exposed by the manual designed augmentations and model training for sequence-level contrastive representation learning, while adopting MSM for the joint-level feature modeling.
-Besides, we further propose a novel clip-level contrastive learning method, which significantly boosts the short-term model capacity, along with an effective post-distillation strategy to achieve a more compact representation space.
+Besides, we further propose a novel clip-level contrastive learning method, which significantly boosts the short-term modeling capacity, along with an effective post-distillation strategy to achieve a more compact representation space.
 Finally, extensive experiments under five downstream tasks, not limited to the single action recognition task used in most previous works, demonstrate the promising generalization capacity of the proposed method.
@@
 An effective SSL schema is proposed, which integrates contrastive learning and MSM to jointly model the representations of different granularity, remarkably benefiting different downstream tasks.
-  • We present a comprehensive benchmark of existing skeleton SSL works as well as a summary of the popular datasets and downstream tasks as well as the insightful analysis from the perspective of model backbones, pre-training paradigms, and future possible directions.
+  • We present the first multi-task benchmark of existing skeleton SSL works with the insightful analysis from the perspective of model backbones, pre-training paradigms, and future possible directions.
 Finally, we demonstrate promising performance of our proposed method on five downstream tasks for versatile action representation learning.
@@
 Then, we comprehensively benchmark existing methods in Sec.
-4, and verify the effectiveness of our proposed approach.
+4 on multiple downstream tasks, and verify the effectiveness of our proposed approach.
 Finally, we conclude and summarize with possible future directions in Sec.
@@
 Generally, a two-stage paradigm is utilized in skeleton-based SSL, i.e., pre-training on pretext tasks first and then fine-tuning on downstream tasks.
-In the pre-training stage, different pretext tasks are designed for deep neural networks, capturing training signals derived from the data itself, called the process of self-supervision [66].
+In the pre-training stage, different pretext tasks are designed for deep neural networks, capturing training signals derived from the data itself, called the process of self-supervision [shurrab2022self].
 After that, the learned knowledge as feature representations is transferred to downstream tasks as shown in Fig.
@@
 Note that some downstream tasks, e.g., motion prediction and 2D-to-3D lifting, are not considered in this survey, because they essentially do not rely on the supervisory signal of human annotation and can serve as pre-training pretext tasks themselves, leading to possible unfair comparison.
-For reviews on these topics, we direct readers to [50, 45].
+For reviews on these topics, we direct readers to [lyu20223d, liu2022recent].
 Next, we first introduce skeleton data as well as its collection in Sec.
@@
 2.1 Human Skeleton Representation
-Skeleton represents a human motion as a sequence of the body keypoint coordinates over time.
-The collection methods of skeleton can be divided into two main categories, i.e., marker-based and markerless methods, shown in Fig.
-1 (a).
+The collection methods of skeleton can be divided into two main categories, i.e., marker-based and markerless methods, shown in Fig. 1 (a).
 Marker-based motion capture (Mocap) systems often rely on inertial measurement units (IMU) or reflective markers, placed on the body to track the movement of humans.
-It can provide reliable skeleton data with even sub-millimeter accuracy [4].
+It can provide reliable skeleton data with even sub-millimeter accuracy [buckley2019role].
 However, this method is too costly in many application scenarios and requires highly trained personnel to operate.
 Meanwhile, it suffers from great inconvenience, e.g., the time-consuming placement process and a requirement for a controlled environment.
-The datasets [31, 75] for generation tasks, e.g., pose estimation, are usually collected in this way to guarantee the accuracy.
+The datasets [ionescu2013human3, von2018recovering] for generation tasks, e.g., pose estimation, are usually collected in this way to guarantee the accuracy.
 Markerless Mocap often depends on deep learning algorithms for pose estimation from RGB and depth data.
 Video RGB data can be easily obtained from the Internet while we can also utilize the camera hardware, e.g., the depth camera including Microsoft Azure Kinect, and single or multiple RGB video cameras, to collect the RGB, infrared, and depth images in deployment.
-Then to get the 3D motion, computer vision algorithms, on multi-view geometry and pose estimation, i.e., OpenPose [6] are then employed to detect and extract joint center locations.
+Then to get the 3D motion, computer vision algorithms, on multi-view geometry and pose estimation, i.e., OpenPose [cao2017realtime] are then employed to detect and extract joint center locations.
 The whole pipeline is presented in Fig.
 1 (a).
-However, due to the limitations of hardware and estimation algorithms, such method can have large errors compared to marker-based methods [15].
-Nevertheless, it is still chosen for most skeleton-based datasets [62, 44, 43] for its simpleness and convenience.
+However, due to the limitations of hardware and estimation algorithms, such method can have large errors compared to marker-based methods [dolatabadi2016concurrent].
+Nevertheless, it is still chosen for most skeleton-based datasets [shahroudy2016ntu, liu2019ntu, liu2020pku] for its simpleness and convenience.
 2.2 SSL Methods for Skeleton
+SSL has been widely explored in the previous literature for the image domain.
+Some early works adopt the context-driven pretext tasks to model the inherent data structure, e.g., predicting the rotation label [gidaris2018unsupervised], correlation map [li2023correlational], and clusters based on the pseudo-labels [caron2018deep, zhan2020online].
+Besides, contrastive-style methods push the performance of SSL to new heights, including invariance decorrelation [zbontar2021barlow, bardes2021vicreg], distillation-based [grill2020bootstrap, chen2021exploring, caron2021emerging, oquab2023dinov2, darcet2023vision], clustering-based [caron2020unsupervised, li2020prototypical].
+Meanwhile, the generative learning, especially the recent masked image modeling [xie2022simmim, he2022masked], has drawn a surge of interest and been proven effective for downstream vision downstream tasks.
+These works have played an important role and motivate the new efforts for skeleton-based SSL.
+However, due to the huge difference between human skeleton and image data as discussed in Sec.
+1, it is suboptimal to transfer these methods directly for skeleton representation learning.
+Therefore, we present this literature review with a high-level methodological taxonomy, and also shed light on the challenges and skeleton-specific designs to motivate further work.
 With respect to the pretext tasks, most existing skeleton-based representation learning methods encompass three categories: (1) context-based, (2) generative learning, and (3) contrastive-learning methods.
@@
 The model is encouraged to learn the spatial-temporal relationships by training on the pre-defined task.
-Emerging from the image domain, the pretext tasks rely on the context understanding, e.g., rotation prediction [21].
+Emerging from the image domain, the pretext tasks rely on the context understanding, e.g., rotation prediction [gidaris2018unsupervised].
 In contrast, skeleton data introduces an additional temporal dimension, and possesses more compact spatial information, which presents a new challenge on how to mine the meaningful spatial-temporal context of skeleton by pretext tasks.
@@
 Due to the variance of the observation viewpoint, the skeleton estimation can suffer from occlusion and noise.
-Therefore, learning view-invariant representations should be beneficial for action recognition task, and has been widely studied in the supervised skeleton-based action recognition task [19].
+Therefore, learning view-invariant representations should be beneficial for action recognition task, and has been widely studied in the supervised skeleton-based action recognition task [gao2022global].
 Li et al.
-[37] proposed a view classification pretext task for unsupervised action representation learning.
+[li2018unsupervised] proposed a view classification pretext task for unsupervised action representation learning.
 Specifically, as shown in Fig.
 3 (a), skeleton sequences of different views are fed into the encoder, subsequent to which a view classifier predicts the view labels.
-To learn the view-invariant features, a Gradient Reversal Layer [18] is added to reverse the optimization direction of the encoder.
+To learn the view-invariant features, a Gradient Reversal Layer [ganin2015unsupervised] is added to reverse the optimization direction of the encoder.
 Resort to this, the encoder can learn the features insensitive to view in an adversarial manner.
 Likewise, Paoletti et al.
-[55] adopted the rotation prediction, i.e., pitch angle, yaw angle, and roll angle, as well as a gradient reverse operation to achieve the viewpoint-invariance learning.
-There are also some subsequent SSL works [87, 53, 3] focusing on learning view-invariant representations.
+[paoletti2022unsupervised] adopted the rotation prediction, i.e., pitch angle, yaw angle, and roll angle, as well as a gradient reverse operation to achieve the viewpoint-invariance learning.
+There are also some subsequent SSL works [yang2022via, men2023focalized, bian2022view] focusing on learning view-invariant representations.
 However, they adopted a contrastive learning paradigm to explicitly learn the alignment between the two views and showed a significant performance improvement.
@@
 3 (b), e.g., with a Cross-Entropy loss.
-Some skeleton-based SSL works [40, 83] are equipped with temporal sorting to model the temporal dependencies, similar to SSL works in video [82, 34].
+Some skeleton-based SSL works [lin2020ms2l, xu2021prototypical] are equipped with temporal sorting to model the temporal dependencies, similar to SSL works in video [xu2019self, lee2017unsupervised].
 However, such pretext task is often integrated with other pretext tasks, because it only models temporal features explicitly, leaving the crucial spatial representations of skeleton under-explored.
@@
 Cheng et al.
-[12] designed a movement direction prediction task.
+[cheng2021hierarchical] designed a movement direction prediction task.
 The model is constrained to estimate the direction of the instantaneous joint velocity, i.e., whether the targeted joints are moving in a positive direction.
 Based on this, Kim et al.
-[33] proposed a multi-interval pose displacement strategy to encourage both local and global attention learning.
+[kim2022global] proposed a multi-interval pose displacement strategy to encourage both local and global attention learning.
 The key idea is to predict the motion direction and magnitude of the central joint and other joints during a random interval.
 On the other hand, Yang et al.
-[91] introduced more motion prior information as the pseudo label, i.e., intra-joint motion variance, inter-joint motion covariance, intra-frame joint angle, and inter-frame motion deviation.
+[yang2022motion] introduced more motion prior information as the pseudo label, i.e., intra-joint motion variance, inter-joint motion covariance, intra-frame joint angle, and inter-frame motion deviation.
 The model is pre-trained by a regression task to predict this prior knowledge.
@@
 It naturally creates an information bottleneck to achieve dimension reduction, mapping data from input space onto low-dimensional feature space.
-Due to its simpleness, AE is used to learn prior representations in earlier skeleton-based works [68, 84].
+Due to its simpleness, AE is used to learn prior representations in earlier skeleton-based works [su2020predict, xu2021unsupervised].
 However, since the skeleton data is still redundant in spatial and temporal dimensions, the encoder tends to memorize the input data at a low level instead of modeling the high-level semantic knowledge.
-Then, denoising AE (DAE) is studied for alleviating this shortcut [69, 54].
+Then, denoising AE (DAE) is studied for alleviating this shortcut [su2021self, nie2020unsupervised].
 For example, Nie et al.
-[54] applied a series of view corruptions and constrained the model to reconstruct the clean data to disentangle the view and pose features.
+[nie2020unsupervised] applied a series of view corruptions and constrained the model to reconstruct the clean data to disentangle the view and pose features.
 On the other hand, inspired by the remarkable success of masked language/image modeling (MLM, MIM), which can be regarded as a variant of DAE, researchers have aroused a surge of interest in exploring masked skeleton modeling (MSM) for human action representation learning.
@@
 However, lacking good scalability, these models only show mediocre performance improvement.
-Recently, inspired by the success of Vision Transformer (ViT) as masked autoencoders (MAE) [25], Transformer-based models have been explored for the masked skeleton modeling [9, 51].
+Recently, inspired by the success of Vision Transformer (ViT) as masked autoencoders (MAE) [he2022masked], Transformer-based models have been explored for the masked skeleton modeling [chen2022hierarchically, mao2023masked].
 However, due to the relative redundancy in skeleton and the lack of a large-scale dataset as ImageNet, directly applying Transformer for masked skeleton reconstruction can suffer from the over-fitting problem.
@@
 (1) for MIM.
-In the field of RGB images, MAE [25] adopts a patch-based masking strategy with a large masking ratio of 75%.
+In the field of RGB images, MAE [he2022masked] adopts a patch-based masking strategy with a large masking ratio of 75%.
 For the skeleton data, researchers have fully explored the masking strategy in the spatial-temporal dimension.
-Although the optimal masking strategy and ratio can differ with the encoder backbone and MSM task setting, most works have found that spatially body-part-based and temporally segment-based strategy with large masking ratio, e.g., 90% in [52], can produce decent results.
+Although the optimal masking strategy and ratio can differ with the encoder backbone and MSM task setting, most works have found that spatially body-part-based and temporally segment-based strategy with large masking ratio, e.g., 90% in [mao2022cmd], can produce decent results.
 Concretely, body-part level masking refers to performing masking regarding the different joints in a body part as a whole, e.g., torso and left leg, while the temporal segment-based strategy masks the same joints across consecutive frames.
-These designs aim to reduce the shortcuts caused by spatial-temporal redundancy in skeleton, which is in line with the observation in images [26] and videos [74].
+These designs aim to reduce the shortcuts caused by spatial-temporal redundancy in skeleton, which is in line with the observation in images [he2022masked] and videos [tong2022videomae].
 Recently, more elaborate masking strategies have been studied to achieve more effective exploitation of valuable semantic information.
-The work [85] found that masking the limbs is always better than masking the torso and head, especially the right hand and leg.
-MAMP [51] proposes a motion-aware masking strategy.
+The work [yan2023skeletonmae] found that masking the limbs is always better than masking the torso and head, especially the right hand and leg.
+MAMP [mao2023masked] proposes a motion-aware masking strategy.
 The moving parts are located and masked by calculating the motion displacement between adjacent frames, which achieves better results than the random masking strategy.
 These results demonstrate that masking motion regions, which are often semantic-rich, promotes the model to learn more meaningful features in masked motion modeling.
-It can be also explained by the theory in [99], i.e., because samples from the same action often contain similar motion patterns, this masking strategy on the motion regions implicitly leads to better alignment of mask-induced positive pairs, achieving more discriminative feature spaces.
+It can be also explained by the theory in [zhang2022mask], i.e., because samples from the same action often contain similar motion patterns, this masking strategy on the motion regions implicitly leads to better alignment of mask-induced positive pairs, achieving more discriminative feature spaces.
 Meanwhile, some literature adopts special mask designs, yielding new MSM task forms.
-MS2[40] directly utilizes temporal frame-level masks to train the model on the motion prediction task.
-MotionBERT [103] adopts a 2D-to-3D lifting pretext task, in which it masks all the values in depth $z$ channel and partially in $x$ and $y$ channels, encouraging the model to predict the original 3D skeletons.
+MS2[lin2020ms2l] directly utilizes temporal frame-level masks to train the model on the motion prediction task.
+MotionBERT [zhu2023motionbert] adopts a 2D-to-3D lifting pretext task, in which it masks all the values in depth $z$ channel and partially in $x$ and $y$ channels, encouraging the model to predict the original 3D skeletons.
 These works can be divided into masked skeleton modeling in a broad sense.
-2) Reconstruction Target, known as the implementation of $\mathcal{T}_{tar}$, also varies from the input space to the feature space.
+2) Reconstruction Target, known as the $\mathcal{T}_{tar}$, also varies from the input space [he2022masked, xie2022simmim] to the feature space [assran2023self, gao2024mimic, xie2022masked, haghighat2023pre, bardes2024revisiting] as the image domain.
 Most existing works reconstruct the skeleton data in the input coordinate space using MSE loss.
-Some works [78] propose to perform temporally reverse reconstruction to learn the skeleton dynamics instead of trivial representations that just remember the input.
+Some works [wang2022contrast] propose to perform temporally reverse reconstruction to learn the skeleton dynamics instead of trivial representations that just remember the input.
 Instead of directly reconstructing the input, Mao et al.
-[51] proposed to reconstruct the motion in MSM, i.e., the difference of the corresponding joints between adjacent frames.
-The work [90] formulates the skeleton as an unordered 3D point cloud and maps the 3D data onto color space.
+[mao2023masked] proposed to reconstruct the motion in MSM, i.e., the difference of the corresponding joints between adjacent frames.
+The work [yang2023self] formulates the skeleton as an unordered 3D point cloud and maps the 3D data onto color space.
 This mapping function is artificially defined according to the spatial-temporal relationship of skeleton joints.
@@
 Generally, contrastive learning pursues the consistency of the positive samples, which are usually the augmented counterparts of the original data.
-MoCo [26] utilizes the negative samples to establish an instance discrimination pretext task.
-Meanwhile, self-distillation [22] and feature decoupling [93, 2] concepts have also been explored in contrastive learning.
-These early pioneering works [26, 8, 22, 93], have made a huge impact and encouraged unique designs for skeleton contrastive learning.
-For skeleton contrastive representation learning, most existing methods are based on MoCo v2 [26].
+MoCo [he2020momentum] utilizes the negative samples to establish an instance discrimination pretext task.
+Meanwhile, self-distillation [grill2020bootstrap] and feature decoupling [zbontar2021barlow, bardes2021vicreg] concepts have also been explored in contrastive learning.
+These early pioneering works [he2020momentum, chen2020simple, grill2020bootstrap, zbontar2021barlow, robinson2020contrastive], have made a huge impact and encouraged unique designs for skeleton contrastive learning.
+For skeleton contrastive representation learning, most existing methods are based on MoCo v2 [he2020momentum].
 It adopts an asynchronously momentum-updated key encoder and an online query encoder, along with a memory queue to store a large number of consistent negative samples.
@@
 $\mathcal{L}=-\log\frac{\exp(z_{q}\cdot z_{k}/\tau)}{\exp(z_{q}\cdot z_{k}/\tau)+\sum_{i=1}{\exp(z_{q}\cdot m_{i}/\tau)}},$ (2)
-where the $z_{q}$/$z_{k}$ is the query/key embedding encoded by the query/key model and $\tau$ is the temperature hyper-parameter.
+where the $z_{q}$$z_{k}$ is the query/key embedding encoded by the query/key model and $\tau$ is the temperature hyper-parameter.
 $m_{i}$ is the $i$-th feature anchor as the negative sample.
@@
 1) Augmentation Design and Learning Strategy.
-Data augmentation exposes novel motion patterns and generates diverse positive views, which have been found crucial to the success of contrastive learning [73, 24].
-Different from images, data augmentation for skeleton is relatively less developed.
-Therefore, the earlier works mainly focus on exploring practical spatial-temporal augmentations for skeleton contrastive learning [60, 72, 20].
+Data augmentation exposes novel motion patterns and generates diverse positive views, which have been found crucial to the success of contrastive learning [tian2020makes, guo2022aimclr].
+Unlike images, data augmentation for skeleton is relatively less developed.
+Therefore, the earlier works mainly focus on exploring practical spatial-temporal augmentations for skeleton contrastive learning [rao2021augmented, thoker2021skeleton, gao2021efficient].
 For example, Rao et al.
-[60] proposed a series of augmentations including Rotation, Shear, Reverse, Gaussian Noise, Gaussian Blur, Joint Mask and Channel Mask, some of which are used as the default basic augmentations in the future research.
+[rao2021augmented] proposed a series of augmentations including Rotation, Shear, Reverse, Gaussian Noise, Gaussian Blur, Joint Mask and Channel Mask, some of which are used as the default basic augmentations in the future research.
 Notably, the optimal augmentations for different backbones are usually different.
-For example, the Joint Mask is found detrimental for GCNs [97], while beneficial for GRU model [72], which implies that GCNs are more sensitive to the spatial corruption.
+For example, the Joint Mask is found detrimental for GCNs [zhang2022hiclr], while beneficial for GRU model [thoker2021skeleton], which implies that GCNs are more sensitive to the spatial corruption.
 Generally, this difference arises on account of different modeling mechanisms and model capacities, which further increases the difficulty of the augmentation design and selection.
 Inspired by the success of mixing-based augmentations for images, Chen et al.
-[94] proposed SkeleMixCLR equipped with SkeleMix augmentation which combines the topological information of different skeleton sequences.
+[chen2022skelemixclr] proposed SkeleMixCLR equipped with SkeleMix augmentation which combines the topological information of different skeleton sequences.
 To learn from this mixed skeleton, SkeleMixCLR obtains the corresponding part level and the whole-body level features and pursues the local-global invariance.
-Based on this, SkeAttnCLR [29] further designs an attention mechanism to perform local contrastive learning on salient and non-salient features.
+Based on this, SkeAttnCLR [hua2023part] further designs an attention mechanism to perform local contrastive learning on salient and non-salient features.
 Due to the generated novel input views and regularization effect on the feature space, mixing augmentation often leads to consistent improvement for representations.
 To further improve consistency learning, researchers have made endeavors to introduce more and stronger data augmentations.
-However, the ensuing problem is the over-distortion [1] of the augmented data, leading to the model performance degradation.
+However, the ensuing problem is the over-distortion [bai2022directional] of the augmented data, leading to the model performance degradation.
 In other words, some strong augmentations would seriously corrupt the semantic information, change the data distribution, and result in the difficulty of model consistency learning.
-To address this, AimCLR [24] utilizes two branches to encode the weakly and strongly augmented views, respectively.
+To address this, AimCLR [guo2022aimclr] utilizes two branches to encode the weakly and strongly augmented views, respectively.
 Then the model optimizes the similarity of the distributions output by two branches as a soft consistency learning target:
-$\mathcal{L}_{Soft}$ $=KL(p(z|z_{weak})\|p(z|z_{strong})),$ (3)
+$\mathcal{L}_{Soft}(z)$ $=KL(p(z|z_{weak})\|p(z|z_{strong})),$ (3)
 ${p}\left(z|z_{*}\right)$ $=\frac{\exp(z\cdot z_{*}/\tau)}{\exp(z_{k}\cdot z_{*}/\tau)+\sum_{i=1}^{M}{\exp(m_{i}\cdot z_{*}/\tau)}},$
@@
 Further, Zhang et al.
-[97] introduced more strong augmentations, e.g., randomly dropping the skeleton edges and joints, and proposed a hierarchical contrastive learning framework.
+[zhang2022hiclr] introduced more strong augmentations, e.g., randomly dropping the skeleton edges and joints, and proposed a hierarchical contrastive learning framework.
 It performs a decoupled progressive augmentation invariance learning by optimizing the consistency only between the augmented samples with adjacent strength, where the weakly augmented branch serves as the mimic target of the strongly augmented branch.
@@
 In addition, Lin et al.
-[41] explicitly distinguished the static regions and motion regions, namely, actionlet, in human skeleton, and introduced a motion-aware augmentation strategy.
+[lin2023actionlet] explicitly distinguished the static regions and motion regions, namely, actionlet, in human skeleton, and introduced a motion-aware augmentation strategy.
 By mining the actionlet in an unsupervised manner, the semantic-reserving augmentations are employed for actionlet regions, while the noise perturbations for non-actionlet regions, avoiding the over-distortion problem.
-On the other hand, instead of augmenting at the input level, HaLP [61] proposes a latent positive hallucinating method by exploring the latent space around the corresponding prototype.
+On the other hand, instead of augmenting at the input level, HaLP [shah2023halp] proposes a latent positive hallucinating method by exploring the latent space around the corresponding prototype.
 In summary, the development of data augmentations along with the corresponding learning strategy significantly boosts the performance of skeleton contrastive learning, which has always been an important and popular topic for 3D skeleton contrastive learning.
@@
 Concretely, false negatives refer to the negative samples but from the same semantic category.
-Traditional contrastive learning relies on the one-hot label and directly pushes them away in the feature space, which forces the model to discard the shared semantic information and leads to slow convergence [30].
+Traditional contrastive learning relies on the one-hot label and directly pushes them away in the feature space, which forces the model to discard the shared semantic information and leads to slow convergence [huynh2022boosting].
 We point out that, due to the lack of description of objects and backgrounds, there are fewer action categories represented by skeleton data, which leads to a more serious FN problem.
 To this end, the following aspects are considered to tackle this issue.
-The first straightforward solution is based on the false negative detection and cancellation [38, 24, 96].
+The first straightforward solution is based on the false negative detection and cancellation [chen2021incremental, huynh2022boosting, li20213d, guo2022aimclr, zhang2022cpm].
 These methods first calculate the similarity between the sample and negatives, selecting the top $k$ negatives that are most similar as the potential false negatives, which are then involved as extended positives in contrastive learning.
-Some other methods [52, 98] suggest employing the distillation objective as an adaptive re-weighting regularization for the one-hot instance discrimination pretext task, which can be formulated as follows:
-$\mathcal{L}_{KD}$ $=-{p}\left({z}_{k},\tau_{k}\right)\log{p}\left(z_{q},\tau_{q}\right),$ (4)
+Some other methods [xie2022delving, mao2022cmd, zhang2023prompted] suggest employing the distillation objective as an adaptive re-weighting regularization for the one-hot instance discrimination pretext task, which can be formulated as follows:
+$\mathcal{L}_{KD}(z_{q},z_{k})$ $=-{p}\left({z}_{k},\tau_{k}\right)\log{p}\left(z_{q},\tau_{q}\right),$ (4)
 ${p_{j}}\left({z},\tau\right)$ $=\frac{\exp({z}\cdot{m}_{j}/\tau)}{\sum_{i=1}{\exp({z}\cdot{m}_{i}/\tau)}}.$
@@
 Specifically, if a sample possesses high similarity with the positive, a larger attraction weight would be assigned to involve it in similarity optimization.
-Besides, prototype-based contrastive learning is also studied for high-level semantic consistency learning.
+Besides, prototype-based contrastive learning is also studied for high-level semantic consistency learning [li2020prototypical, caron2020unsupervised, xu2021prototypical].
 It usually performs clustering in feature space to assign instances to different cluster prototypes as a pseudo-semantic label.
 Then the model learns more high-level semantics by contrasting different prototypes.
-By virtue of such a way, the model pays more attention to the cluster-level discrimination task rather than the instance-level, which alleviates the false negative problem.
+By virtue of such a way, the model pays more attention to the cluster-level discrimination task rather than instance-level to alleviate the false negative problem.
 In addition to the above designs, as introduced above, some negative-sample-free contrastive frameworks have been developed recently, which avoid the difficulty of explicitly specifying negative examples.
-Based on BYOL [22], the works [17, 7] performs the positive-only consistency learning in the hyperbolic space.
-Some other works [95, 70], inspired by the feature decorrelation concept, propose to learn the decorrelated representations based on Barlow twins [93] and Variance-Invariance-Covariance Regularization (VICReg) [2] as the framework.
+Based on BYOL [grill2020bootstrap], the works [francohyperbolic, chen2023self] performs the positive-only consistency learning in the hyperbolic space.
+Some other works [zhang2022skeletal, sun2023unified], inspired by the feature decorrelation concept, propose to learn the decorrelated representations based on Barlow twins [zbontar2021barlow] and Variance-Invariance-Covariance Regularization (VICReg) [bardes2021vicreg] as the framework.
 Figure 4: Different representations of skeleton data. From left to right are time series, 2D pseudo-image, spatial-temporal graph.
+Figure 5: Summary of the advantages/limitations of different skeleton-based SSL methodologies.
 3) Cross-Modal Learning.
-Skeleton naturally provides different data modalities to represent human movement, e.g., motion and bone modality defined in the previous works [63, 10].
-Meanwhile, many works [63, 10] have proved that the fusion of different modal knowledge can be beneficial to achieve richer and more representative action depictions.
-Therefore, cross-modal learning has attracted the attention of researchers for skeleton representation learning.
-CRRL [78] performs contrastive learning between the joint and motion modalities, where the encoders for motion and joint data are homologous, i.e., obtained by momentum-updated strategy.
+Skeleton naturally provides different data modalities to represent human movement, e.g., motion and bone modality defined in the previous works [shi2019two, chen2021channel].
+Meanwhile, many works [shi2019two, chen2021channel] have proved that the fusion of different modal knowledge can be beneficial to achieve richer and more representative action depictions.
+For example, joint contains essential information about human movements while lacking the explicit encoding of some high-order information.
+Bone view can capture the structural patterns well.
+However, it is subject to the predefined human model and still lacks temporal information.
+Motion explicitly encodes the temporal dynamics while losing the spatial information.
+It is promising to explore these different modalities to achieve better action representation.
+CRRL [wang2022contrast] performs contrastive learning between the joint and motion modalities, where the encoders for motion and joint data are homologous, i.e., obtained by momentum-updated strategy.
 However, this design can be unreasonable because the model can have difficulty dealing with two modalities simultaneously.
-In contrast, CrosSCLR [38] adopts three separate encoders for joint, bone, and motion data, and aims to align the distributions among the neighborhoods in latent space.
+In contrast, CrosSCLR [li20213d] adopts three separate encoders for joint, bone, and motion data, and aims to align the distributions among the neighborhoods in latent space.
 Intuitively, it encourages the sample embedding together with its neighbors of modality $v$ should also be close in the latent space of modality $u$.
 However, this leads to a two-stage training process, as the model needs to be pre-trained on a single view first to obtain reliable unique modal knowledge.
-To this end, CMD [52] proposes a general bidirectional distillation objective, where the modal knowledge is directly modeled as the whole similarity distribution in the customized latent space.
+To this end, CMD [mao2022cmd] proposes a general bidirectional distillation objective, where the modal knowledge is directly modeled as the whole similarity distribution in the customized latent space.
 The proposed cross-modal distillation objective and the instance discrimination task jointly optimize the model, yielding a concise single training phase.
-On the other hand, to get rid of the limitation of using separate encoders for different modalities, UmURL [70] develops a unified modality-agnostic encoder, which can handle different modal inputs relying on the modality-specific input embedding layer and feature projection layer.
+On the other hand, to get rid of the limitation of using separate encoders for different modalities, UmURL [sun2023unified] develops a unified modality-agnostic encoder, which can handle different modal inputs relying on the modality-specific input embedding layer and feature projection layer.
 Meanwhile, we can also organize skeleton data into different representations as shown in Fig.
 4, e.g., the temporal series, 2D pseudo-images with frame and joint dimensions, and spatial-temporal graphs with nodes and edges.
-Following this concept, ISC [72] encodes these different skeleton representations with different model backbones, which are then projected into a shared latent space.
+Following this concept, ISC [thoker2021skeleton] encodes these different skeleton representations with different model backbones, which are then projected into a shared latent space.
 A cross-modal contrastive learning loss is applied for pre-training.
@@
 In addition to the aforementioned literature, some works also employ contrastive learning.
-Considering the hierarchical structure of skeleton data, HiCo [16] performs contrastive learning in a hierarchical manner to model features of different levels.
-Some works [40, 83, 71, 78] combine the contrastive learning paradigm with others in a multi-tasking manner.
-Notably, in addition to simply combining reconstruction and contrastive learning pretext tasks, PCM3 [98] proposes a collaborative design to further improve the representation learning.
-Eq-Contrast [42] formulates the temporal sorting task into an equivariant contrastive learning objective for multi-task pre-training.
-Meanwhile, not limited to SSL, contrastive learning is also studied in supervised [101] and semi-supervised [81, 65] learning for skeleton.
+Considering the hierarchical structure of skeleton data, HiCo [hico2023] performs contrastive learning in a hierarchical manner to model features of different levels.
+Some works [lin2020ms2l, xu2021prototypical, tanfous2022and, wang2022contrast] combine the contrastive learning paradigm with others in a multi-tasking manner.
+Notably, in addition to simply combining reconstruction and contrastive learning pretext tasks, PCM3 [zhang2023prompted] proposes a collaborative design to further improve the representation learning.
+Eq-Contrast [lin2024mutual] formulates the temporal sorting task into an equivariant contrastive learning objective for multi-task pre-training.
+Meanwhile, not limited to SSL, contrastive learning is also studied in supervised [zhou2023learning] and semi-supervised [xu2022x, shu2022multi] learning for skeleton.
 2.3 Summary and Discussion
@@
 All these methods aims to capture joint features as well as their relationships from unlabeled skeleton data, to obtain a meaningful representation of the motion.
-Due to the lack of ground-truth labels, these methodologies introduce prior knowledge in different means to enable representation learning.
-Specifically, context-based methods introduce the prior knowledge by the predefined pretext-task designs, e.g., predicting the equivariant properties of some specific transformation, learning the spatial concurrence or temporal dependencies of skeleton joints.
-Generative approaches generally reconstruct or predict the skeleton target without much prior information.
-However, some recent works have shown that some special designs, e.g., masking the motion regions, can further boost the performance, where the prior knowledge is also utilized that the moving parts can better represent the action.
-In the context of contrastive learning, the prior information is mostly reflected through the data augmentations for the construction of positive pairs, which affect the invariance learning of transformation-related features.
-Overall, although downstream tasks are often unknowable in pre-training, these methods can obtain a potentially beneficial representation space by introducing different prior knowledge.
+Due to the lack of ground-truth labels, these methodologies introduce prior knowledge in different means to enable representation learning, e.g., context task design, augmentation choices, and masking strategy.
+To summary the different skeleton-based SSL methods based on our taxonomy, we present the advantages and limitations of each methodology category in Fig.
+5.
+Context-based methods are often simple to implement, flexible for design, and training-friendly compared with the other two paradigms.
+However, as evidenced by the previous literature, the main concern is the limited model capacity of generalization and semantic understanding due to the gap between the low-level pretext tasks and the downstream tasks.
+Contrastive learning is recognized by its highly discriminative representations with strong performance, especially under linear evaluation protocol.
+More importantly, it has been proven effective for different backbones for skeleton modeling.
+The other side of the coin is augmentation sensitivity, false negative problem, and high training cost as previously discussed.
+Meanwhile, some researchers point out the concern of dimensional collapse in contrastive learning [jingunderstanding, lin2024mutual], where the embedding vectors only span a lower dimensional space, limiting the representation capacity.
+Finally, the masked skeleton modeling is discussed considering its dominant position in generative learning.
+It is well compatible with Transformer, with expected strong generalization performance, especially under fine-tuning protocol.
+However, performance can significantly drop when employing other backbones, which can limit its deployment in some low-latency, resource-constrained scenarios.
+Meanwhile, it is also concerned about masking strategy sensitivity and time-consuming training.
+To end this section and possibly motivate future work, we further summarize and emphasize the special design of skeleton-based SSL approaches that go beyond the high-level SSL methodology taxonomy.
+$\bullet$ Context Pretext Tasks Based on Human Topology and Kinematics.
+Different from images, skeleton pretext tasks focus on the spatial-temporal data structure learning of human body, e.g., temporal sorting, angle, and displacement prediction.
+$\bullet$ Data Augmentation for Novel Motion Patterns.
+Skeleton augmentation has always been the focus to boost contrastive learning.
+These efforts include the spatial and temporal, invariant and equivariant [lin2024mutual], weak and strong [guo2022aimclr, zhang2022hiclr], intra and inter augmentations [chen2022skelemixclr, zhang2023prompted] as well as the corresponding learning strategy.
+$\bullet$ Cross-Modal Knowledge Modeling.
+Skeleton naturally provides different data modalities (or views) to represent human movement, e.g., joint, motion, and bone, enabling cross-modal knowledge modeling for better action representations.
+$\bullet$ False-Negative Cancellation.
+As discussed, the diversity of human motions is much less than natural images, leading to a more serious false negative problem, which is an important topic in skeleton contrastive learning.
+$\bullet$ Masking Strategy in MSM. As a crucial component in MSM, spatial-temporal masking strategy has also been widely studied specifically for the highly structured human skeleton data.
+$\bullet$ Multi-Granularity Representation Learning.
+Skeleton data naturally presents a hierarchy, e.g., sequence-level, clip-level, and joint-level.
+It can be beneficial to capture these action clues at different scales, enriching the action representation and alleviating the over-fitting problem.
 3 The Proposed Method
 3.1 Motivation
-By reviewing the previous works, we find that most existing skeleton-based SSL methods employ a single paradigm or simply combine different methods, evaluated on the action recognition task solely.
-We argue this ignores the hierarchical design for representation learning of different granularity and leaves the generalization capacity of model leashed, as we will verify later in Sec.
-4.7.
-To construct a strong baseline model for skeleton-based SSL, we propose prompted contrast with masked motion modeling (PCM3++), which combines contrastive learning and MSM by exploring the potential collaboration.
+By reviewing the previous works, we find that most existing skeleton SSL methods employ a single paradigm or simply combine different methods with action recognition task for evaluation solely, which limits the model capacity of learning representations at different granularity and levels and leaves the generalization capacity leashed.
+To this end, we propose prompted contrast with masked motion modeling (PCM3++), which combines contrastive learning and MSM by exploring potential collaboration.
 It enables versatile representation learning by joint, clip, and sequence level feature modeling, and significantly improves the performance of different downstream tasks.
-Specifically, our method is based on our previous work [98] and integrates valuable improvement for better versatile representation learning.
-We first briefly introduce the joint-level and sequence-level feature modeling of the baseline in Sec.
-3.2 - 3.4.
-Based on this, we introduce a novel asymmetric clip-level contrastive learning method in Sec.
-3.2.
-The training strategy is presented in Sec.
-3.5, including the newly proposed post-distillation policy for further representation refinement.
-The readers can refer to [98] for more details.
+Table 1: Main notations and the definition.
+Notation Definition
+$s^{\prime}(s_{ach})$, $z^{\prime}(z_{ach})$ The global key-branch intra-augmented
+skeleton and its embedding.
+$s_{intra}$, $z_{intra}$ The query-branch intra-augmented
+skeleton and its embedding.
+$s_{inter}$, $z_{inter}$ The query-branch inter-augmented
+skeleton and its embedding.
+$s_{clip}$, $z_{clip}$ The query-branch temporal-augmented
+skeleton clip and its embedding.
+$s_{mask}$, $z_{mask}$ The query-branch masked
+skeleton and its embedding.
+$s_{predict}$, $z_{predict}$ The reconstructed skeleton from $s_{mask}$
+and its embedding.
+$z_{inter}^{\prime}$ The linear interpolation of two skeleton
+embeddings in mix inter-augmentation.
+$p_{*}$ The prompts of different views.
+$f(\cdot),h(\cdot),dec(\cdot)$ The encoder, projector, and decoder.
+Specifically, our method is based on our previous work [zhang2023prompted], which combines contrastive learning and masked skeleton modeling in a synergetic manner.
+Based on this, we further extend and strengthen the design with the key idea, i.e., versatile action representation learning.
+In the following, we first introduce the contrastive learning (Sec.
+3.2) and MSM (Sec.
+3.3) scheme involved in baseline method as well as their collaboration design (Sec.
+3.4).
+Based on this, we introduce our new improvements in PCM3++ compared to [zhang2023prompted].
+Concretely, an effective clip-level contrastive learning method is designed in Sec.
+3.2.(3), which significantly boosts the capacity of short-term motion modeling as illustrated in Fig.
+6.
+Besides, a post-distillation strategy is developed in Sec.
+3.5.(2) to further boost the representation compactness.
+These advancements further improve the representation generalization for various downstream tasks.
+For clarity, the main notations are summarized in Table 1.
 3.2 Skeleton Contrastive Learning
-Our pipeline for contrastive learning follows MoCo V2 [26].
+Our pipeline for contrastive learning follows MoCo V2 [he2020momentum].
 For a positive pair $(z,z^{\prime})$, the model optimizes the InfoNCE objective defined in Eq.
@@
 $\centering\mathcal{L}_{CL}(z,z^{\prime})=\mathcal{L}_{Info}(z,z^{\prime})+\mathcal{L}_{KD}(z,z^{\prime}).\@add@centering$ (5)
+The intuition behind this is to combine the one-hot hard labels in InfoNCE to explicitly cluster features while avoiding collapse, and the soft labels in the distillation objective, i.e., the similarity scores, to achieve fine-grained relational knowledge learning, which also alleviates the false negative problem as discussed in Sec.
+2.2.3.(1).
 To boost the consistency learning, we propose a series of effective spatial-temporal augmentations to construct diverse positive pairs, which are introduced as follows.
+Figure 6: The overview of the proposed method.
+We integrate masked skeleton modeling and contrastive learning paradigms to learn the joint- (blue part), clip- (yellow part), and sequence-level (green part) representations.
+Meanwhile, a series of domain-specific prompts are introduced as training guidance.
 1) Spatial Intra- and Inter- Skeleton Contrastive Learning.
@@
 Subsequently, we can obtain the embeddings corresponding to the mixed data by $z_{inter}=h_{q}\circ f_{q}(s_{inter})$.
-The optimized objective is for $(z_{inter},z_{inter}^{\prime})$, where $z_{inter}^{\prime}=(1-\lambda)(h_{k}\circ f_{k}(s_{1}))+\lambda(h_{k}\circ f_{k}(s_{2}))$.
+The optimized objective is $\mathcal{L}_{CL}(z_{inter},z_{inter}^{\prime})$ in Eq.
+(5), where $z_{inter}^{\prime}=(1-\lambda)(h_{k}\circ f_{k}(s_{1}))+\lambda(h_{k}\circ f_{k}(s_{2}))$.
 2) Temporal Asymmetric Clip Contrastive Learning.
-The above designs focus on the sequence-level consistency modeling, i.e., the augmented data often contains sufficient temporal motion information.
-However, it ignores the short-term motion representation learning of model, which is necessitated for downstream tasks with dense prediction such as action detection.
-Meanwhile, the consistency learning for short clips, which can be viewed as a challenging temporal augmentation, can also benefit the whole representation quality learned by model.
-To this end, we present the clip-level skeleton contrastive learning design in this part.
-To sample a motion sub-sequence (clip), we can apply the Temporal Crop transformation, and the obtained clip is much shorter ($10\%\sim 40\%$) than the original sequence.
+The above designs only consider the sequence-level consistency modeling, i.e., the augmented data often contains sufficient temporal motion information.
+However, it ignores the short-term motion modeling, which is necessitated for downstream tasks of dense prediction such as action detection.
+On the other hand, the consistency learning for short-term motions, e.g., clips, which can be viewed as challenging temporal augmented views, can also benefit the whole representation quality learned by model.
+To this end, we present the clip-level contrastive learning design.
+To sample a motion clip, we can apply the Temporal Crop transformation, and the obtained clip is much shorter ($10\%\sim 40\%$) than the original sequence.
 However, the clip only contains partial motion information and directly aligning the semantic embeddings of two clips is difficult for model, leading to the unstable training.
-Therefore, we introduce an asymmetric design, which only feeds the short clip into query branch, while the key branch takes the normal original sequence as an anchor.
+Therefore, we introduce an asymmetric design, which only feeds the short clip into query branch, while the key branch takes the pre-crop original sequence as an anchor.
 Meanwhile, we keep the two augmented data possessing the same spatial views to decouple the temporal variation.
@@
 It means that, $s_{ach}$ is subsuming $s_{clip}$ in temporal dimension while sharing exactly the same spatial transformations, to provide a precise and feasible target for clip-level consistency learning.
-Similarly, the model takes the positive pair $(z_{clip},z_{ach})$ for optimization, where $z_{clip}=h_{q}\circ f_{q}(s_{clip})$.
+Similarly, the model takes the positive pair $(z_{clip},z_{ach})$ for $\mathcal{L}_{CL}$ optimization in Eq.
+(5), where $z_{clip}=h_{q}\circ f_{q}(s_{clip})$.
 In implementation, we just mask $z_{ach}=z^{\prime}$ for efficiency.
@@
 The MSE loss between original data $s$ and predicted data $s_{predict}$ by decoder is optimized in the masked regions:
-$\mathcal{L}_{Mask}=\mathbb{E}\left(||(s-dec\circ f_{q}(s_{mask}))\odot(\mathbf{1}-M)||_{2}\right),$ (6)
+[…]
 where $M$ is the binary mask and $\mathbf{1}$ is an all-one matrix with the same shape as $M$.
 3.4 On the Connection of Contrastive Learning and Masked Prediction
-As discussed in the work [98], simply combining these two tasks can be sub-optimal due to the inherent gap of their feature modeling mechanisms [57, 23].
+As discussed in the work [zhang2023prompted], simply combining these two tasks can be sub-optimal due to the inherent gap of their feature modeling mechanisms [qi2023contrast, gui2023survey].
 To this end, we explore the potential synergy for exploitation between them.
@@
 $\{(z_{q},z_{k})\}=\{($ $z_{intra},z^{\prime}),(z_{inter},z^{\prime}_{inter}),(z_{clip},z^{\prime}),$ (7)
-(((( $z_{mask},z^{\prime}),(z_{predict},z^{\prime})\}.$
+$($ $z_{mask},z^{\prime}),(z_{predict},z^{\prime})\}.$
+They are fed into the encoders to optimize the $\mathcal{L}_{Info}$ and $\mathcal{L}_{KD}$ in Eq. 5.
 2) High-Level Semantic Guidance.
@@
 (6), leading to better masked prediction learning and higher quality of $s_{predict}$ as positive samples.
+With these synergetic designs, the masked skeleton modeling provides novel positives to the contrastive learning, boosting the model generalization capacity.
+On the other hand, the supervision signals as gradients of contrastive learning branch also assists the masked prediction branch with the high-level semantic guidance provided by inter-sample contrast.
+These designs alleviate the gap of feature modeling mechanism of these two tasks, which limits the performance of previous works that directly combine them, and hence yield better representation quality.
 3.5 The Whole Training Strategy
@@
 $\mathcal{L}_{Info}^{all}=\sum_{(z_{q},z_{k})}\mathcal{L}_{Info}^{query},\ \ \mathcal{L}_{KD}^{all}=\sum_{(z_{q},z_{k})}\mathcal{L}_{KD}^{query}.$
-$\mathcal{L}_{Info}^{query}$ and $\mathcal{L}_{KD}^{query}$ are the respective terms for the specific $query$ view summarized in Eq.
+$\mathcal{L}_{Info}^{query}$ and $\mathcal{L}_{KD}^{query}$ are the respective terms for the specific queryquery view summarized in Eq.
 (7).
@@
 After obtaining a good prior feature space by Eq.
-(8), we directly remove the one-hot label constraint in InfoNCE objective $\mathcal{L}_{Info}^{all}$, and only apply the soft label, i.e., the distillation loss $\mathcal{L}_{KD}^{all}$, which assigns the attraction weights adaptively according to the similarity.
+(8), we directly remove the one-hot label constraint in InfoNCE objective $\mathcal{L}_{Info}^{all}$, i.e.,
+$\mathcal{L}=\lambda_{m}\mathcal{L}_{Mask}+\lambda_{kd}\mathcal{L}_{KD}^{all},$ (10)
+where we only apply distillation term that assigns the attraction weights adaptively according to the similarity.
 This can be seen as a more explicit feature clustering process to obtain a more compact representation space by alleviating the false negative problem.
+Note this single distillation objective would not cause model collapse due to the well initialized representation space learned in previous pre-training stage.
+3.6 Discussion
+We summarize our method as a direct response to the challenges mentioned in previous literature review.
+We target at the versatile action representation learning, adopting our previous work [zhang2023prompted] as the baseline for its effectiveness to learn multi-granularity features.
+Building on this, we further integrate the clip-level contrastive learning with it to present a more comprehensive solution with joint, clip, and sequence level modeling.
+As a response to previously discussed challenges, 1) we carefully design the spatial-temporal augmentations involving intra- and inter- data samples to present a strong contrastive learning baseline, which eases the efforts of subsequent works.
+2) Further, the proposed multi-granularity representation modeling can effectively enhance the feature diversity, alleviating the dimensional collapse implicitly according to [zhang2022mask].
+3) Meanwhile, for the false negative problem, we employ a different perspective, i.e., post-training to deal with it.
+In the following sections, we will present the quantitative analysis on these designs.
 4 Benchmark Skeleton SSL Methods
@@
 Meanwhile, the evaluation of proposed PCM3++ is reported.
-Table 1: Summary of popular skeleton datasets. * denotes the collected by Marker-MoCap methods.
-Dataset Instance Class Sub. View Joint
-CMU* [13] 2235 45 144 - 31
-SBU [92] 300 8 7 - 15
-NW-UCLA [76] 1,494 10 10 3 20
-UWA3D [58] 1,075 30 10 5 15
-SYSU [28] 480 12 40 1 20
-NTU 60 [62] 56,880 60 40 80 25
-PKUMMD [43] 21,545 51 66 3 25
-Kinetics [32] $\sim$260k 400 - - 18
-TSU [14] 16,115 31 18 - 17
-NTU 120 [44] 114,480 120 106 155 25
-UAV [39] 22,476 155 119 17
-BABEL* [56] 48,978 120 - - 25
-4.1 Datasets
-We first summarize the popular datasets for skeleton action understanding in Table 1. Among them, the following datasets are widely used in skeleton-based SSL evaluation.
-1) NTU RGB+D 60 Dataset (NTU 60) [62] is the most popular skeleton dataset.
+We first introduce the following datasets, which are widely used in skeleton-based SSL evaluation. The summary of more datasets for skeleton action understanding can be found in Supp.
+1) NTU RGB+D 60 Dataset (NTU 60) [shahroudy2016ntu] is the most popular skeleton dataset.
 There are 56,578 videos with 25 joints for a human, captured by three Microsoft Kinect v2 cameras.
@@
 b) Cross-View (xview): the training set consists of front and two side views of the action performers, while testing set includes the left and right 45 degree views.
-2) NTU RGB+D 120 Dataset (NTU 120) [44] is an extension to NTU 60 dataset.
+2) NTU RGB+D 120 Dataset (NTU 120) [liu2019ntu] is an extension to NTU 60 dataset.
 There are 114,480 videos collected with 120 action categories, performed by 106 subjects.
@@
 b) Cross-Setup (xset): the training data uses even setup IDs, while testing data are odd setup IDs.
-3) PKU Multi-Modality Dataset (PKUMMD) [43] is another large-scale benchmark with available skeleton data.
+3) PKU Multi-Modality Dataset (PKUMMD) [liu2020pku] is another large-scale benchmark with available skeleton data.
 Two subsets, Part I and Part II, are provided.
@@
 Convolutional neural networks (CNNs) are utilized to process the skeleton sequence as a pseudo-2D image.
-Meanwhile, skeletons can also be treated as point clouds, and hence the Dynamic Graph CNN (DGCNN) [79] is also explored as the feature extractor.
+Meanwhile, skeletons can also be treated as point clouds, and hence the Dynamic Graph CNN (DGCNN) [wang2019dynamic] is also explored as the feature extractor.
 4.2 Downstream Tasks
@@
 It can also be referred to as the Temporal Action Localization or Action Segmentation.
-Following the previous works [49, 43], the Average Precision (mAP) at different temporal Intersection over Union (tIoU) thresholds between the predicted and the ground truth intervals is utilized as the metric.
+Following the previous works [luo2018graph, liu2020pku], the Average Precision (mAP) at different temporal Intersection over Union (tIoU) thresholds between the predicted and the ground truth intervals is utilized as the metric.
 4) Occluded Action Recognition focuses on the action understanding with occlusion, which is prevalent in human activity.
@@
 4.3 Implementations Details
-For the proposed method, we follow the implementation and experiment settings of previous works [98, 52].
+For the proposed method, we follow the implementation and experiment settings of previous works [zhang2023prompted, mao2022cmd].
 For the existing methods, we give priority to the reported results in their original paper.
@@
 Details can be found in Supp.
-4.5 Benchmark on Action Recognition Task
-We benchmark different SSL methods on NTU-60 and NTU-120 benchmarks using the widely adopted action recognition task in Table LABEL:tab:main_comp.
-More comprehensive comparison on additional datasets can be found in the supplementary material.
-The feature dimension of the last layer before the classifier is provided, which often correlates with the upper capacity of the model.
-For the evaluation of other tasks, the community still lacks unified benchmarks.
-To this end, we reproduce some popular works with the same settings as our work to obtain comparable results in Sec.
-4.6.
-1) Comparison across Different Pretext Tasks.
+4.4 A Multi-Task Benchmark for Skeleton-based SSL
+In this section, we present our multi-task benchmark for skeleton-based SSL methods, which is the first time to the best of our knowledge, to benefit the community.
+Considering most previous SSL works are evaluated only on action recognition, we split the benchmark into two parts for presentation, i.e., a full comparison on action recognition task and the reproduction benchmark on other tasks.
+First, we curate a comprehensive comparison of the widely adopted action recognition task in Table LABEL:tab:main_comp with the backbone and pretext task category mark.
+On the other hand, for other different tasks, an available benchmark is provided covering action retrieval, occluded recognition, action detection, and few-shot learning tasks.
+We select the recent popular works with official code released, and make huge efforts to reproduce them on the same settings to present a multi-task comparison as shown in Table LABEL:tab:other_tasks.
+4.4.1 Comparison on Action Recognition Task
+1) Across Pretext Tasks.
 As shown in Fig.
-5, context-based and generative learning pretext tasks are mainly studied in the earlier works first.
+7, context-based and generative learning pretext tasks are mainly studied in the earlier works.
 However, these methods usually fail to achieve a satisfactory performance under linear evaluation due to the modeling of too many low-level features and the lack of effective design.
-Recently, MAMP [51] adopts MSM with large mask ratios and motion-aware reconstruction targets, obtaining a highly discriminative semantic feature space for both linear evaluation and fully fine-tuning protocol, demonstrating promising results for Transformer-based generative learning.
+Recently, MAMP [mao2023masked] adopts MSM with large mask ratios and motion-aware reconstruction targets, obtaining a highly discriminative semantic feature space for both linear evaluation and fully fine-tuning protocol.
+The subsequent works [lin2025idempotent, wu2025macdiff] also demonstrate promising results for Transformer-based MSM.
+However, these methods are only reported based on the joint view, which makes cross-modal knowledge ensemble in this paradigm still an open problem.
 Contrastive learning has always been a popular topic triggered by the success in the image field, e.g., SimCLR and MoCo.
 The model can learn a separable high-level representation space, making it dominant in learning linear representations.
-These works start from the exploiting the baseline algorithms [60, 72], and significantly boost the performance by applying stronger augmentations [24, 97, 41], effective training strategies [96, 7], cross-modal knowledge [38, 52], achieving rapid improvement in recent three years.
-Besides, combining different pre-training paradigms [98, 42] for skeleton-based representation learning has demonstrated promising results recently.
+These works start from the exploiting the baseline algorithms [rao2021augmented, thoker2021skeleton], and significantly boost the performance by applying stronger augmentations [guo2022aimclr, zhang2022hiclr, lin2023actionlet], effective training strategies [zhang2022cpm, chen2023self], cross-modal knowledge [li20213d, mao2022cmd], achieving rapid improvement in recent three years.
+Besides, combining different pre-training paradigms [zhang2023prompted, lin2024mutual] for skeleton-based representation learning has demonstrated promising results recently.
 This indicates that the representations learned by different pre-training paradigms can be complementary and beneficial.
-2) Comparison across Different Model Backbones.
+2) Across Model Backbones.
 Different model backbones often employ different pre-training tasks.
@@
 7 in SSL, and the Transformer models are also promising recently.
-For fully fine-tuning protocol, Transformer models are more popular due to the strong representation power.
-Remarkably, with self-supervised pre-training, a vanilla Transformer [52] has surpassed the GCN/Transformer with complex designs in supervised training [35, 64], demonstrating the significant role for alleviating the over-fitting problem and generalization capacity improvement of SSL.
-[…]
-Method xview xsub
-1% 10% 1% 10%
-Semi-Supervised Methods
-X-CAR [81] - 78.2 - 76.1
-MAC-Learning [65] - 78.5 - 74.2
-Self-Supervised Methods
-LongT GAN [100] - - 35.2 62.0
-MS2L [40] - - 33.1 65.1
-ISC [72] 38.1 72.5 35.7 65.9
-CMD [52] 53.0 80.2 50.6 75.4
-HiCo [16] 54.8 78.3 54.4 73.0
-PCM3 [98] 53.1 82.8 53.8 77.1
-PCM3++ 57.5 84.5 57.1 79.4
-[…]
-4.6 Benchmark on Different Downstream Tasks
-To give a thorough evaluation of generalization capacity of current SSL models and our method PCM3++, we conduct extensive experiments and perform benchmarking on five downstream tasks, including action recognition, action retrieval, occluded recognition, action detection, and few-shot learning.
-For comparison, we choose the popular SSL methods mostly based on GRU model, which achieves a strong performance currently, while also ensuring a fair comparison with our GRU-based approach.
-1) Skeleton-based Action Recognition. The following two evaluation settings are adopted.
-Firstly, the unsupervised learning results using linear evaluation protocol is shown in Table LABEL:tab:main_comp, and PCM3++ achieves better or competitive results compared with other state-of-the-art methods on different evaluation protocols.
-Meanwhile, the results on PKUMMD dataset can be found in supplementary material, which also indicates the strong generalization capacity on the noisy data.
-In addition, we also report the semi-supervised setting results, which fine-tunes the whole model/classifier with only a portion of labeled data.
-As shown in Table 2, PCM3++ obtains better performance than both the semi- and self-supervised methods with different proportions of training data, and only fine-tuning the classifier tends to be better with extremely little data.
-In contrast, previous works based on reconstruction [100] or contrastive learning [72, 52] solely can not achieve satisfactory results.
-Table 3: The results of action retrieval with joint stream.
-Method NTU 60 (%) NTU 120 (%)
-xsub xview xsub xset
-ISC [72] 62.5 82.6 50.6 52.3
-HiCLR [97] 60.6 73.1 46.0 46.0
-CRRL [78] 60.7 75.2 - -
-CMD [52] 70.6 85.4 58.3 60.9
-HaLP [61] 65.8 83.6 55.8 59.0
-HiCo [16] 67.9 84.4 55.9 58.7
-MAMP [51] 62.0 70.0 - -
-UmURL [70] 71.3 88.3 58.5 60.9
-PCM3 [98] 73.7 88.8 63.1 66.8
-PCM3++ 75.4 89.4 64.5 67.1
-Table 4: The results of occluded action recognition.
-Method Occluded NTU 60
-Spatial Occ. Temporal Occ.
-xsub xview xsub xview
-MoCo-GRU [26] 64.8 72.6 68.8 74.8
-ISC [72] 62.8 70.6 68.9 76.8
-CRRL [78] 56.8 61.4 61.0 66.2
-CMD [52] 67.1 72.7 72.7 79.5
-HiCo [16] 66.4 72.1 71.2 76.7
-PCM3 [98] 80.8 87.0 77.6 86.1
-PCM3++ 81.8 88.0 79.9 85.8
-2) Skeleton-based Action Retrieval.
-Following previous work [68], a K-nearest neighbors (KNN) classifier ($k$=1) is adopted to retrieve the nearest training sample for each testing data in the representation space.
-The precision is reported as accuracy in Table 3.
-Our method achieves a competitive performance compared with other latest methods.
-This indicates the highly distinguishable representations learned by our method.
-Meanwhile, it is found that contrastive learning based methods usually perform better than reconstruction-based methods, e.g., LongT GAN [100] and MAMP [51].
-Table 5: Action detection results on PKUMMD Part I benchmark under linear evaluation protocol.
-Method mAP@tIoU (%)
-0.1 0.3 0.5
-Supervise 63.4 61.4 53.4
-Unsupervised Methods
-MoCo-GRU [26] 68.2 67.2 63.4
-CRRL [78] 57.6 55.7 52.1
-ISC [72] 64.6 62.9 58.7
-CMD [52] 73.7 72.8 68.4
-HiCo [16] 51.8 50.8 45.8
-LAC [88] 55.2 - -
-PCM3 [98] 73.3 72.8 68.2
-PCM3++ 75.5 74.6 69.8
-Table 6: Comparison of few shot skeleton-based action recognition. ($k$, $n$) represents $k$-way $n$-shot.
-Method xset (%) xsub (%)
-(5, 1) (5, 5) (5, 1) (5, 5)
-Supervise 64.5 81.2 67.3 85.2
-Unsupervised Methods
-MoCo-GRU [26] 58.5 78.7 61.1 81.3
-CRRL [78] 58.3 77.1 55.2 75.0
-ISC [72] 62.3 81.3 64.4 83.6
-CMD [52] 64.9 82.8 68.6 85.7
-HiCo [16] 64.6 81.7 67.1 84.2
-PCM3 [98] 65.2 82.9 68.8 86.8
-PCM3++ 66.7 84.8 69.9 87.3
-3) Action Recognition with Occlusion.
+For fully fine-tuning protocol, GRU model is less effective due to the over-fitting caused by high dimension and huge parameters.
+Instead, Transformer models are more popular due to the powerful attention mechanism.
+Remarkably, with self-supervised pre-training, a vanilla Transformer [mao2023masked] has surpassed the GCN/Transformer with complex designs in supervised training [lee2023hierarchically, shi2020decoupled], demonstrating the significant role for alleviating the over-fitting problem and generalization capacity improvement of SSL.
+4.4.2 Comparison on Other Downstream Tasks
+Table 2: Analytics on the pre-training of different datasets for multiple downstream tasks.
+Pre-training Recognition Retrieval Occluded Detection Few Shot
+NTU 60 NTU 120 NTU 60 NTU 120 NTU 60 xview PKUMMD NTU 120 xset
+xsub xsub xsub xsub Spatial Temporal mAP@0.1 mAP@0.5 (5,1) (5,5)
+NTU 60 84.8 73.5 75.4 58.8 88.0 85.8 75.5 69.8 66.7 84.8
+NTU 120 85.1 76.7 75.3 64.5 87.7 88.2 75.3 69.4 69.6 87.2
+PKUMMD 77.6 67.5 65.1 52.9 78.0 79.3 70.7 65.7 58.9 78.9
+NTU 120+PKUMMD 85.8 77.0 76.2 64.5 88.4 88.0 76.2 70.5 69.4 87.0
+1) Experiment Setting. We introduce the implementations of each task first and more details can be found in Supp.
+Skeleton-based Action Retrieval.
+Following previous work [su2020predict], a K-nearest neighbors (KNN) classifier (kk=1) is adopted to retrieve the nearest training sample for each testing data in the representation space.
+All models are first pre-trained and then evaluated on the target dataset.
+The precision is reported as accuracy.
+Action Recognition with Occlusion.
 We evaluate the transfer ability of representations learned from the clean dataset to the action recognition with occluded data.
 We adopt the linear evaluation protocol.
-Following [67], we construct a synthetic occluded dataset on both spatial and temporal dimension.
+Following [song2019richly], we construct a synthetic occluded dataset on both spatial and temporal dimension.
 For spatial dimension, different body parts are randomly masked.
@@
 The testing set is constructed by the same masks across different methods, with a masking ratio of [0.3, 0.7].
-As presented in Table 4, our method can capture underlying structures in the distorted data owing to the masked contrastive learning, and is good at dealing with the spatial occlusion.
-Meanwhile, this capacity extends well for temporal occlusion, demonstrating a desirable generalization capacity.
-However, due to the lack of modeling learning of occluded data, other methods are with poor robustness for occlusion.
-4) Skeleton-based Action Detection.
-We follow the previous works [49, 43], and evaluate the short-term motion modeling capacity by action detection task.
+Skeleton-based Action Detection.
+We follow the previous works [luo2018graph, liu2020pku], and evaluate the short-term motion modeling capacity by action detection task.
 We train the attached linear classifier to predict frame-level categories to produce the final proposal.
 The encoder is pre-trained on NTU 60 dataset and then we fix it, transferring the learned representations to untrimmed PKUMMD Part I dataset.
-The mean average precision (mAP) of different actions is adopted as metric with different temporal overlapping ratios (tIoU) in Table 5.
-We provide a fully fine-tuned model (train both the encoder and classifier) without pre-training as the baseline.
-First, it can be found the SSL pre-training can significantly improve the performance compared with the baseline model.
-Besides, our method integrates the clip-level representation contrastive learning and achieves the best scores compared with latest CMD [52], HiCo [16], and LAC [88], which ignore the clip-level representation modeling.
+The mAP of different actions is adopted as metric with different tIoUs.
+Unsupervised Few-shot Learning.
+We evaluate the performance of SSL as few-shot learners follow the previous work [lu2022self].
+Specifically, the model is first pre-trained on NTU 60 dataset.
+Then, a simple classifier, e.g., Support Vector Machine (SVM), is fitted on the output features by the pre-trained encoder of the support data set.
+Finally, the adapted classifier along with the encoder is utilized to infer the query samples.
+We select 20 new categories in NTU 120111Specifically, the classes (index from 0) 60, 61, 66, 69, 72, 78, 79, 80, 84, 90, 91, 95, 96, 98, 99, 100, 102, 106, 108, 111, 113 114, 115 are selected.
+which are not seen in pre-training, as the support/query sets.
+2) Comparison and Analytics. The results are summarized in Table LABEL:tab:other_tasks. We utilize the officially released model weights, keeping the experiments as fair and aligned as possible.
+We first present the discussion on the perspective of model backbones.
+GRU models usually employ higher representation dimensions due to the smaller GPU memory consumption, which makes a more informative representation available, leading to stronger performance in retrieval task than GCNs and Transformers.
+GCNs model the skeleton more compactly with structured prior knowledge, resulting in a better capacity to handle over-fitting, accounting for the stronger few-shot performance for unseen categories.
+However, the other side of the coin is the sensitivity to occlusion, where GCN can suffer from serious representation degradation because of the distribution shift caused by occlusion.
+Besides, it is found the GRU-based often performs better than the GCN model in terms of action detection, which primarily focuses on the spatial structure while ignoring the fine-grained temporal dependency.
+In contrast, GRUs model the skeletons in a temporal-based auto-regressive manner and possess better scalability for the long-sequences and short-clips.
+Besides, the Transformer-based model [mao2023masked] also demonstrates promising results when employing the masked skeleton modeling pretext task, and can be expected with a better fine-tuning performance due to the powerful attention mechanism.
+On the taxonomy of pretext tasks, due to the lack of promising open-sourced context-based methods, we mainly focus on contrastive and generative learning methods.
+First, it is found that contrastive learning methods usually perform better than reconstruction-based methods, e.g., MAMP [mao2023masked] and CRRL [qi2023contrast], in action retrieval and few-shot learning.
+These tasks mainly rely on the direct measure in the representation space, i.e., cosine similarity, indicating the dominant position of contrastive learning for achieving highly compact representations.
+Remarkably, the masked modeling method, e.g., MAMP [mao2023masked], achieves significant results in action detection task.
+We believe that masked modeling with Transformer architecture is very promising for handling the action detection task.
+By splitting the long sequences into short clips as patches for masked prediction, the learned representations can well capture fine-grained action cues and achieve notable results.
+However, it is not observed with as good performance under occluded recognition as expected, although it explicitly involves masked data modeling, which can be explored in future work.
+4.4.3 Cross-Dataset Analytics for Different Downstream Tasks
+We study the effect of the dependency relationship between different pre-training datasets on multiple downstream tasks with the proposed method.
+Specifically, we pre-train the model on NTU 60, NTU 120, PKUMMD, and their hybrid dataset, respectively, which is then evaluated on various downstream tasks under the transfer learning setting.
+To achieve a better understanding, we first introduce the relationship between different datasets in terms of the data composition.
+Concretely, NTU 120 extends NTU 60 on the action categories from 60 to 120 without introducing new samples of the original categories.
+PKUMMD is another skeleton dataset with similar collection settings, sharing a significant overlap of category distribution with NTU 60 and possessing fewer samples, which can be approximated as a smaller version of NTU 60.
+The quantitative results are presented in Table 2.
+$\bullet$ First, as expected, reducing the scale of pre-training data generally results in performance degradation for downstream tasks, especially for the evaluation of the large-scale NTU 120 dataset.
+This is because the new classes in testing are unseen during pre-training, resulting in limited generalization capacity.
+[…]
+$\bullet$ In contrast, for the tasks requiring to explicitly generalize to new data distributions, e.g., temporal occluded recognition with distribution shift by impairing data structure (Note only spatial occlusion is seen in the pre-training) and few shot tasks with new categories, introducing new categories can benefit the performance significantly.
+This indicates more essential knowledge of human motion is enabled for model to boost generalization.
+To sum up, we analyze the effect of pre-training on different datasets to downstream tasks from two factors, category distribution and the sample number within each class, providing valuable insights for the skeleton SSL dataset choice and construction.
+Overall, a better performance can be expected with a larger pre-training dataset.
+However, it is also necessary to consider the actual downstream scenarios and the adopted pre-training methods to adjust data component.
+Table 3: Ablation study on the novel positive pair and semantic guidance.
+Novel Positive Pair Semantic Guidance xview
+$s_{mask}$ $s_{predict}$
+88.2
+89.3
+90.0
+90.5
 Table 4: Ablation study on the clip-level contrastive learning and the post-distillation.
@@
 w post-distillation 84.8 91.0 75.5
-5) Unsupervised Few-shot Learning.
-In this task, we evaluate the performance of SSL as few-shot learners follow the previous work [48].
-Specifically, the model is first pre-trained on NTU 60 dataset.
-Then, a simple classifier, e.g., Support Vector Machine (SVM) in our implementation, is fitted on the output features by the pre-trained encoder of the support data set.
-Finally, the adapted classifier along with the encoder is utilized to infer the query samples.
-We select 20 new categories in NTU 120111Specifically, the classes (index from 0) 60, 61, 66, 69, 72, 78, 79, 80, 84, 90, 91, 95, 96, 98, 99, 100, 102, 106, 108, 111, 113 114, 115 are selected.
-which are not seen in pre-training, as the support and query sets following [46].
-In Table 6, our method generalizes well on the unseen categories and surpasses the supervised baseline and other methods notably, although there is no training for unseen categories.
-6) Compared with the Baseline Method.
-We evaluate and present the comparison results of our baseline model, PCM3 [98], on various downstream tasks as shown in Table LABEL:tab:main_comp-6.
+4.5 Analysis of PCM3++
+1) Evaluating the PCM3++.
+As shown in the previous results in Table LABEL:tab:main_comp and Table LABEL:tab:other_tasks, the proposed method achieves remarkable results on different downstream tasks compared with the state-of-the-art methods, and demonstrates significant improvement over our baseline method [zhang2023prompted].
 As our new improvement, we integrate an effective clip-level contrastive learning scheme and present a novel post-distillation training strategy.
-These methodological advancements consistently improve the model’s representation quality and robustness on various downstream tasks, especially the semi-supervised action recognition and action detection performance.
-To sum up, self-supervised learning provides a strong solution to boost various downstream tasks for versatile skeleton-based action understanding.
-Generally, a higher recognition performance roughly indicates a better representation space for most downstream tasks, as found in Table 3 and Table 6, but not equal to greater robustness to occlusion.
-Meanwhile, action detection is another crucial task for action analysis, which mainly focus on the short-term motion modeling.
-We design a novel asymmetric clip-level contrastive learning pretext task and effectively boost the detection performance.
-Remarkably, our proposed method achieves promising results on these tasks compared with other methods that ignore the versatile representation learning of different granularity.
-4.7 Analysis of PCM3++
-1) Analysis of clip contrastive learning.
+These methodological advancement is studied as follows.
+2) Analysis of clip contrastive learning.
 We add the temporal clips as the asymmetric positive sample of the anchor sequence, which further improves the representation quality as shown in Table 4.
@@
 Our designs boost versatile representation learning of different granularity and the effect on more downstream tasks can be found in Supp.
-2) Effect of the post-distillation design.
+3) Effect of the post-distillation design.
 As shown in Table 4, the post-distillation training strategy can bring further improvement slightly for the recognition task.
 It alleviates the false negative problem by removing the one-hot pseudo label in the contrastive learning, achieving a more compact representation space.
-Table 8: Results of different backbones.
-Method Backbone xsub xview
-ActCLR [41] GCN 80.9 86.7
-CSTCN [77] GRU 83.1 88.7
-HiCLR [97] Transformer 76.6 80.8
-PCM3++ GCN 80.6 85.1
-PCM3++ GRU 84.8 91.0
-PCM3++ Transformer 80.2 84.9
-Table 9: FLOPs and Params results.
-[…]
-GL-Transformer [33] 214M 59.4G 83.8%
-ISC [72] 106M 13.7G 85.2%
-CMD [52] 99M 17.3G 86.9%
-PCM3 [98] 103M 15.0G 90.4%
-PCM3++ 103M 17.9G 91.0%
-3) Results of different backbones.
-We give more results with different backbones in Table 8.
-Specifically, for GCN and Transformer, we adopt ST-GCN [86] and DSTA-Net [64] following previous works [41, 97].
-As we can see, our method shows good generalization capacity across different model backbones, and performs better or on-par compared with the latest SSL methods.
-Additional improvement can be potentially achieved by further searching for hyper-parameters on different backbones, which is, however, not the focus of this paper, and we leave it for future work.
-We finally adopt GRU model in implementation for the more friendly memory usage of GPU and the higher performance.
-4) Complexity Analysis.
+Table 5: FLOPs, params, and training time (all tested on a single NVIDIA A40 GPU) results.
+Model Params FLOPs Training Acc.
+Time
+GL-Trans. [kim2022global] 214M 59.4G 53.62h 83.8%
+ISC [thoker2021skeleton] 106M 13.7G 45.49h 85.2%
+CMD [mao2022cmd] 99M 17.3G 44.13h 86.9%
+PCM3 [zhang2023prompted] 103M 15.0G 45.31h 90.4%
+PCM3++ 103M 17.9G 47.95h 91.0%
+4) Novel positive pair and semantic guidance.
+We analyze the effect of synergetic design between masked prediction and contrastive learning in Table 3.
+As we can see, the masked views and predicted views as positives bring notable improvement, indicating the potential benefits of masked prediction training to contrastive learning.
+On the other hand, we propagate the gradients of contrastive branch to masked prediction as high-level semantic guidance rather than stop the gradient.
+This boosts masked prediction training and yields more informative positive samples, leading to further performance improvement.
+5) Complexity analysis.
 We give an analysis of space and computational complexities of our method for pre-training in Table 5.
-Compared with other GRU-based and Transformer-based methods, our method achieves a significant performance improvement with an acceptable cost of the complexity.
+Compared with other GRU-based and Transformer-based methods, our method achieves a significant performance improvement with an acceptable cost of the complexity and training time.
 For the space complexity, the main extra cost is the reconstruction decoder (4M of space).
@@
   • Multi-Modal Learning.
-The exploration of multi-modal pre-training including human skeleton data is still insufficient, e.g., text, infrared data, and RGB images.
+The exploration of multi-modal pre-training including human skeleton data is still insufficient, e.g., texts, and RGB images.
 For example, as a useful complement to skeleton, RGB data can provide additional background knowledge that can boost action representation learning.
+Some recent works [petrovich2023tmr, yu2024mopatch] explore the human motion-text representation learning.
+However, the insufficiency of high-quality motion-text datasets still limits the model generalization capacity.
   • Towards Skeleton in the Wild.
@@
 As discussed in previous sections, most works only focus on the recognition task, leaving the generation capacity across different tasks of SSL models under-explored.
-Meanwhile, new frameworks can also be explored, e.g., Diffusion model [27], which is promising to handle both generative and discriminative downstream tasks.
+Meanwhile, new frameworks can also be explored, e.g., Diffusion model [ho2020denoising], which is promising to handle both generative and discriminative downstream tasks.
 Declarations
   • Competing interests.
-This work was supported in part by the National Natural Science Foundation of China under Grant No.62172020, and in part by the Key Laboratory of Science, Technology and Standard in Press Industry (Key Laboratory of Intelligent Press Media Technology).
+This work was supported in part by the Beijing Major Science and Technology Project under Contract No.
+Z251100008425023, in part by the National Natural Science Foundation of China under Grant 62471009, and in part by the Key Laboratory of Science, Technology and Standard in Press Industry (Key Laboratory of Intelligent Press Media Technology).
   • Data availability.
-The adopted NTU-RGB+D [62, 44] and PKU-MMD [43] dataset are popular public benchmarks in skeleton-based action understanding.
+The adopted NTU-RGB+D [shahroudy2016ntu, liu2019ntu] and PKU-MMD [liu2020pku] dataset are popular public benchmarks in skeleton-based action understanding.
 The code for data processing can be found in https://github.com/JHang2020/PCM3.